### v1

# ConCISE: Confidence-guided Compression in Step-by-step Efficient Reasoning

Abstract
Large Reasoning Models (LRMs) perform strongly in complex reasoning tasks via Chain-of-Thought (CoT) prompting, but often suffer from verbose outputs caused by redundant content, increasing computational overhead, and degrading user experience. Existing compression methods either operate post-hoc pruning, risking disruption to reasoning coherence, or rely on sampling-based selection, which fails to intervene effectively during generation. In this work, we introduce a confidence-guided perspective to explain the emergence of redundant reflection in LRMs, identifying two key patterns: Confidence Deficit, where the model reconsiders correct steps due to low internal confidence, and Termination Delay, where reasoning continues even after reaching a confident answer. Based on this analysis, we propose ConCISE (Confidence-guided Compression In Step-by-step Efficient Reasoning), a framework that simplifies reasoning chains by reinforcing the model’s confidence during inference, thus preventing the generation of redundant reflection steps. It integrates Confidence Injection to stabilize intermediate steps and Early Stopping to terminate reasoning when confidence is sufficient. Extensive experiments demonstrate that fine-tuning LRMs on ConCISE-generated data yields significantly shorter outputs, reducing length by up to ~50% under SimPO, while maintaining high task accuracy. ConCISE consistently outperforms existing baselines across multiple reasoning benchmarks.
ConCISE: Confidence-guided Compression in Step-by-step Efficient Reasoning
Ziqing Qiao1, Yongheng Deng1, Jiali Zeng2, Dong Wang1, Lai Wei1, Fandong Meng2, Jie Zhou2, Ju Ren1, Yaoxue Zhang1 1Tsinghua University 2Pattern Recognition Center, WeChat AI, Tencent Inc., China
1 Introduction
In recent years, Large Language Models (LLMs) have achieved remarkable progress in natural language understanding and generation, particularly in complex reasoning tasks. Large Reasoning Models (LRMs), such as OpenAI-o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025), which leverage the Chain-of-Thought (CoT) paradigm Wei et al. (2022), have demonstrated state-of-the-art performance on benchmarks involving mathematical problem solving and logical inference Muennighoff et al. (2025); Qwen Team (2024). However, a prominent limitation of LRMs is their tendency to generate excessively verbose reasoning chains Feng et al. (2025); Chen et al. (2024). This inefficiency impairs user experience and incurs considerable computational overhead, posing challenges for real-time applications and deployment in resource-constrained settings Team et al. (2025).
Figure 1: Our proposed method, ConCISE, achieves a more favorable trade-off between reasoning compression and task performance than existing methods.
To mitigate the verbosity of reasoning outputs, recent research has focused on compressing the reasoning chains generated by LRMs Qu et al. (2025); Sui et al. (2025). A mainstream and effective strategy is finetuning LRMs on concise reasoning datasets, enabling the LRMs to generate shorter reasoning chains directly during inference Ma et al. (2025); Chen et al. (2024); Munkhbat et al. (2025). The quality of such compression heavily relies on how these datasets are constructed. One common approach is sampling-based selection, which generates multiple candidate reasoning chains and selects the shortest correct one Team et al. (2025), or removes redundant reflection steps after the final answer is obtained Chen et al. (2024). While simple, these approaches lack control over generation and rely on the reasoning chains naturally generated by the model, which may still include unnecessary steps, reducing compression effectiveness. Another line of work relies on post-hoc pruning, where redundant or less informative reasoning steps are identified and removed from a complete reasoning chain Cui et al. (2025); Xia et al. (2025). However, such modifications can disrupt the continuity of reasoning, leading to degraded performance of LRMs after fine-tuning.
To address the limitations of existing approaches, we seek to construct reasoning chains that are both compact and logically intact, effectively removing redundant reflection steps without sacrificing the performance of LRMs after fine-tuning. To this end, we analyse the reasoning behavior of LRMs and identify that the reflection behavior is not solely determined by correctness, but is heavily influenced by the model’s internal confidence. In particular, insufficient confidence can cause the model to re-evaluate correct steps, resulting in redundant reflections. We thus adopt a confidence-guided perspective to explain why LRMs generate reflection steps. From this view, we identify two key patterns underlying redundant reflection: Confidence Deficit, where the model reconsiders correct intermediate steps due to low internal confidence; and Termination Delay, where reflection continues even after a confident final answer is reached. These patterns substantially inflate reasoning chains and offer actionable insights for constructing concise reasoning data.
Building on this analysis, we propose ConCISE (Confidence-guided Compression In Step-by-step Efficient reasoning)—the first framework that leverages a confidence-guided perspective to actively construct concise reasoning datasets by suppressing redundant reflection during generation. ConCISE consists of two complementary components—Confidence Injection and Early Stopping—designed to mitigate Confidence Deficit and Termination Delay, respectively. Confidence Injection inserts confidence phrases before potential reflection points to strengthen the model’s belief in intermediate steps and suppress unnecessary reflection. Early Stopping employs a lightweight confidence detector to monitor the model’s internal confidence and terminate reasoning once sufficient confidence is achieved. Through the synergy of these two mechanisms, ConCISE enables a more efficient and compact reasoning chain. We then fine-tune LRMs using ConCISE-generated data via SFT Wei et al. (2021) and SimPO Meng et al. (2024), and evaluate their performance on multiple reasoning benchmarks. Experimental results show that ConCISE achieves a superior trade-off between reasoning compression and task performance. As shown in Figure 1, ConCISE reduces average response length by ~50% under SimPO while maintaining high accuracy; under SFT, it also significantly outperforms existing methods in compression efficiency. These results demonstrate that LRMs fine-tuned with ConCISE-generated data learn to avoid generating redundant reflection steps, achieving efficient reasoning without sacrificing performance.
The main contributions of this work are summarized as follows:
  • We present the first confidence-guided explanation of redundant reflection in LRM reasoning, identifying key patterns that underlie reasoning verbosity.
  • We propose ConCISE, a framework that proactively suppresses redundant reflection during inference and constructs high-quality, concise reasoning data for LRM fine-tuning.
  • Through extensive evaluation on multiple reasoning benchmarks, we demonstrate that LRMs fine-tuned with ConCISE-generated data achieve a better trade-off between reasoning chain compression and task performance.
2 Related Work
Figure 2: The ConCISE reasoning pipeline applied to solving the equation $73x-2=$ 7. The figure compares the original 8-step reasoning process with the streamlined 5-step process generated by ConCISE. It illustrates how ConCISE models reflection based on confidence values, identifying Confidence Deficit and Termination Delay, and applying Confidence Injection and Early Stopping to suppress redundant reasoning. Here, $C_{i}$ denotes the internal confidence of each step, and $T_{i}$ represents its corresponding threshold.
Recent research has increasingly focused on mitigating verbosity and redundancy in reasoning chains generated by Large Reasoning Models (LRMs)Qu et al. (2025). While CoT enables step-by-step reasoning, it often produces long outputs that increase latency, computational costs, and may degrade final accuracyWu et al. (2025); Nayab et al. (2024); Wang et al. (2025). To address this, existing approaches can be broadly categorized into three types: input-based methods, decoding-time interventions, and model-based training strategies Sui et al. (2025).
Input-based Methods aim to promote concise reasoning by modifying input prompts without changing the model architecture or decoding strategyLee et al. (2025). Common techniques include imposing token limits or instructing the model to reason briefly Han et al. (2024); Renze and Guven (2024). However, these approaches often fail when the model does not consistently follow the instructions. Another approach, task routing, dynamically selects an appropriate model or reasoning strategy based on input characteristics like question complexity Ong et al. (2024); Chuang et al. (2025); Aytes et al. (2025). Its effectiveness, however, depends on the accurate assessment of input characteristics, which is not always reliable.
Decoding-based Methods intervene during inference to control the reasoning process. One direction compresses intermediate steps into latent representations rather than explicit text, which improves brevity but sacrifices interpretability Hao et al. (2024); Shen et al. (2025b). Another strategy uses dynamic decoding, where lightweight models or heuristics evaluate each reasoning step and decide whether to retain, modify, or discard it during generation Sun et al. (2024); Zhang et al. (2025); Xu et al. (2025). While effective at reducing verbosity, these methods introduce additional computational overhead and may disrupt the reasoning flow.
Model-based Methods train models to generate concise reasoning directly. One approach employs reinforcement learning with reward functions that penalize verbosity Shen et al. (2025a); Aggarwal and Welleck (2025); Luo et al. (2025); Arora and Zanette (2025), but such methods can be unstable, sensitive to reward formulation, and impose significant training costs. More commonly, LRMs are fine-tuned on datasets with concise reasoning chainsMa et al. (2025); Cui et al. (2025); Liu et al. (2024); Xia et al. (2025). These datasets are often created by selecting the shortest correct chains or removing redundant parts post-hoc. Although effective, these methods lack a principled understanding of redundancy, potentially retaining unnecessary steps or removing useful context, which may hurt compression efficiency or model performance.
Motivated by the limitations of existing methods, we propose a model-based framework ConCISE. Specifically, we achieve precise identification and suppression of redundant reflection steps throughout the entire reasoning process by analyzing and explaining reflection behavior from a confidence-guided perspective. This mechanism allows the model to generate concise and efficient reasoning chains while maintaining the flow of reasoning. By doing so, we construct a high-quality training dataset and effectively transfer this capability to the model through fine-tuning, enabling it to perform concise reasoning without compromising model performance.
3 Method
3.1 Confidence-guided Formulation of Reflection in LRM Reasoning
We observe that reflection in LRMs is not solely triggered by correctness judgments. In many cases, reflection steps are invoked even on correct reasoning steps, including those that have already been verified, leading to substantial redundancy. This suggests that reflection behavior is also linked to the model’s internal confidence rather than correctness alone. To explain this phenomenon, we adopt a confidence-guided perspective to formalize when and why LRMs engage in reflection.
Let $S_{i}={s_{1},s_{2},\ldots,s_{i}}$ denote the partial reasoning chain up to step $i$, where each $s_{j}$ is a textual reasoning unit. We associate each step $s_{i}$ with a confidence score $c_{i}\in[0,1]$, representing the model’s internal belief in the correctness of that step. The model’s generation policy $\pi_{\theta}$ maps the current reasoning context $S_{i}$ to the next step $s_{i+1}$.
To model the decision between proceeding and reflecting, we introduce a dynamic threshold $t_{i}\in[0,1]$, which may vary with the model architecture or the evolving context. At each step, the model compares its internal confidence $c_{i}$ with this threshold $t_{i}$ to determine whether to continue reasoning or enter reflection. Formally:
$s_{i+1}=\pi_{\theta}(S_{i})=\text{ReflectionStep}\quad\text{if }c_{i}<t_{i}$
Guided by this formulation, we analyze the reflective behavior of LRMs and formally characterize two recurrent patterns responsible for reflection-related redundancy: Confidence Deficit and Termination Delay. These patterns explain how the model’s internal confidence leads to the occurrence of redundant reflection. We will elaborate on each in the following subsection.
(a) Reflection probability after confidence injection.
(b) Detected confidence progression during reasoning.
Figure 3: Validation experiments for Confidence Injection and Termination Delay. The dataset used in these experiments is described in Appendix B.2.
3.2 Two Key Patterns of Redundancy
Confidence Deficit.
One major source of redundancy in LRMs stems from their tendency to undertrust their correct intermediate steps. LRMs often display unexpected hesitation despite exhibiting fine-grained reasoning capabilities and achieving higher stepwise accuracy, triggering reflection even on simple and unambiguous reasoning steps. This self-undermining behavior leads to redundant reflection with minimal semantic gain, ultimately increasing computational cost and disrupting the efficiency of the reasoning process. We refer to this phenomenon as Confidence Deficit. Formally, Confidence Deficit can be defined as the occurrence of a step $s_{i}$ satisfying:
$\exists i,\quad c_{i}<t_{i}\quad\text{despite }s_{i}\text{ being correct}.$
Termination Delay.
Beyond Confidence Deficit, LRMs exhibit another redundant reasoning pattern we term Termination Delay: After producing a confident final answer, the model is expected to conclude with minimal additional reasoning. However, it often continues to generate unnecessary reflection steps even after repeatedly verifying the same conclusion.
This behavior is attributed to a sharp increase in the confidence threshold $t_{i}$ after reaching the answer at step $i^{\prime}$, making it increasingly difficult for the model’s internal confidence $c_{i}$ to exceed $t_{i}$. As a result, even when $c_{i}$ becomes relatively high, it may still fall short of the heightened $t_{i}$, preventing the model from terminating, leading to verbose post-answer reasoning. Formally, Termination Delay can be characterized as:
$\exists i>i^{\prime},\;\forall j<i^{\prime},\quad t_{i}\gg t_{j}\;\text{and}\;c_{i}<t_{i}$
3.3 The ConCISE Framework
To mitigate the above patterns—Confidence Deficit and Termination Delay—we propose ConCISE, an integrated framework that dynamically steers the reasoning generation process to reduce redundant reflections and produce concise reasoning chains.
ConCISE integrates two core mechanisms: Confidence Injection, which proactively inserts confidence phrases to suppress unnecessary reflection steps during reasoning, and Early Stopping, which halts further generation once sufficient post-answer confidence is detected. Together, these mechanisms operate in synergy to generate concise reasoning chains without compromising inference quality. The overall process is illustrated in Figure 2.
3.3.1 Confidence Injection
To alleviate Confidence Deficit, we design a simple yet effective mechanism called Confidence Injection. The key idea is to boost the model’s internal confidence during reasoning by inserting designed phrases, referred to as confidence phrases, to prevent unnecessary reflection steps. Considering that indiscriminate injection could disrupt reasoning, we selectively insert confidence phrases at critical points where the model’s confidence is relatively low—that is, when $c_{i}<t_{i}$.
Specifically, at each reasoning step $s_{i}$, the model first generates the next step $s_{i+1}=\pi_{\theta}(S_{i})$, where $S_{i}$ denotes the reasoning chain up to step $i$. If $s_{i+1}$ is identified as a reflection step, indicating insufficient confidence, we retroactively modify the input by appending a confidence phrase $p_{i}$ sampled from a curated pool $\mathcal{P}$, and regenerate $s_{i+1}$. Formally, the updated generation process is defined as:
$s_{i+1}=\pi_{\theta}(S_{i})\quad\text{if}\;c_{i}\geq t_{i}\;\text{else}\quad\pi_{\theta}(S_{i},p_{i})$
We manually constructed the initial version of confidence phrase pool $\mathcal{P}$ based on observations of certainty in original reasoning chains. To ensure generalization, the pool includes diverse formulations, and confidence phrases are randomly sampled from it during inference. To evaluate and further refine the confidence phrase pool, we conducted experiments measuring the effectiveness of different phrases. Specifically, we inserted each candidate phrase at reflection positions and recorded the probability that the model still generated a reflection step. This probability indicates whether the updated confidence $c_{i}^{\prime}$ (after phrase insertion) exceeds the dynamic threshold $t_{i}$.
As shown in Figure 3(a), different phrases exhibit varying levels of effectiveness, some notably reduce the likelihood of reflection, while others have only a marginal impact. Notably, even the best-performing phrases still yield a reflection rate above 20%, suggesting that confidence injection reduces redundant reflections while retaining necessary verification. Based on these results, we selected the top 20 phrases with the lowest reflection rates to form the final confidence phrase pool. The composition of the phrase pool and further experimental details are provided in Appendix A.1.
3.3.2 Early Stopping
Although Confidence Injection effectively boosts the model’s confidence during reasoning, it does not fully mitigate the pattern of Termination Delay, which is caused by the significant increase in the dynamic threshold $t_{i}$ after the model generates an answer, leading the model to continue unnecessary reflections even when its confidence is already high. To address this, we design an Early Stopping mechanism based on direct estimation of the model’s internal confidence. Specifically, we construct a lightweight confidence detector to provide a quantitative proxy for the model’s internal confidence at each step after the first answer is generated.
We introduce a probing prompt (such as "So I’m") after the current reasoning context and analyze the probability distribution of generated continuations. Based on statistical analysis, we collect a set of confidence-indicative tokens $\mathcal{W}_{+}$, which reflect affirmations of high certainty. Formally, the detected confidence score $\hat{c}_{i}$ at step $i$ is calculated as:
$\hat{c}_{i}=\sum_{w\in\mathcal{W}_{+}}p(w\mid S_{i},\text{Probing prompt}),$
To validate the confidence detector and examine how model confidence evolves post-FAS(First Answer Step), we calculate the average detected confidence $\hat{c}_{i}$ across key reasoning stages, including FAS and the subsequent five reflection steps (Ref1 to Ref5). As shown in Figure 3(b), the model exhibits low confidence at the FAS, but after the first reflection, confidence increases sharply and continues to rise through subsequent reflections. Notably, the final reflection step has an average detected confidence of $0.93$, suggesting that for $i>i^{\prime}$, the dynamic threshold $t_{i}$ effectively stabilizes around this high value, motivating the introduction of a manually controlled lower threshold. Therefore, we define an early stopping threshold $t_{\text{e}}=0.5$. The model is instructed to terminate generation as soon as its detected confidence $\hat{c}_{i}$ exceeds $t_{\text{e}}$. Further details regarding the selection of $t_{e}$ and the confidence detector are provided in Appendix A.2.
In addition, we also compute the average and maximum $\hat{c}_{i}$ before the FAS (Avg Pre-FAS and Max Pre-FAS), both of which remain close to zero. This indicates that our confidence detector only becomes effective after the first answer is generated, ensuring that premature termination during early reasoning stages does not occur.
3.4 Building Efficient Reasoning Chains
ConCISE integrates Confidence Injection and Early Stopping to dynamically control the reasoning generation process. Given an input question, the LRM generates the reasoning chain step-by-step. At each step, the model first examines whether the next generated step is a reflection step. If a reflection is detected, Confidence Injection is applied by appending a confidence prompt and regenerating the step with reinforced internal confidence. After generating the next reasoning step, a lightweight confidence detector estimates the detected confidence score $\hat{c}_{i}$. If $\hat{c}_{i}$ exceeds a predefined threshold $t_{e}$, the model terminates generation, assuming sufficient confidence in the answer. Otherwise, the generated step is appended to the current reasoning chain, and the process repeats.. Algorithm 1 summarizes the overall coordinated reasoning process in ConCISE.
Algorithm 1 Algorithmic workflow of ConCISE for generating concise reasoning chains.
1:Preparation:
2:    Large Reasoning Model $\pi_{\theta}$
3:    Confidence Phrase Pool $\mathcal{P}$
4:    Early Stopping Threshold $t_{e}=0.5$
5:    Prompt Template $T$
6:Input: Question $Q$
7:Initialize: Reasoning chain $S_{0}=T(Q)$
8:for each reasoning step $i=1,2,\ldots$ do
9:     Generate $s_{i}=\pi_{\theta}(S_{i-1})$
10:     if $s_{i}$ is a reflection step then
11:         Apply Confidence Injection:
12:             Sample a confidence phrase $p_{i}\in\mathcal{P}$
13:             Regenerate $s_{i}=\pi_{\theta}(S_{i-1}+p_{i})$
14:     end if
15:     Compute detected confidence $\hat{c}_{i}$
16:     if $\hat{c}_{i}>t_{e}$ then
17:         Terminate
18:     end if
19:     Accept $s_{i}$ and update $S_{i}=S_{i-1}+s_{i}$
20:end for
3.5 Training Objective and Fine-tuning
To align LRM generation with the behavior encouraged by ConCISE, we first construct a fine-tuning dataset by applying the ConCISE pipeline to a set of reasoning questions. This yields concise reasoning traces in which redundant reflections are actively suppressed and timely termination is encouraged through the use of Confidence Injection and Early Stopping.
We fine-tune the model on this dataset using two standard learning paradigms: supervised fine-tuning (SFT), which directly imitates the revised reasoning chains, and Simple Preference Optimization (SimPO), which further amplifies the model’s preference for confident, efficient reasoning. Both approaches optimize a shared objective that governs the desired generation behavior:
$\begin{cases}\pi_{\theta}(S_{i})\rightarrow\pi_{\theta}(S_{i}+p_{i}),&\text{if}c_{i}<t_{i}\\ \pi_{\theta}(S_{i})\rightarrow\text{Terminate},&\text{if }i\geq i^{\prime}\text{ and }\hat{c}_{i}>t_{e}\end{cases}$
This objective encourages the model to learn a generation policy guided by internal confidence, rather than merely imitating surface text patterns. Specifically, when the model’s confidence $c_{i}$ at a given step falls below the threshold $t_{i}$, it learns to approximate the distribution conditioned on an injected confidence phrase $p_{i}$, which helps improve internal certainty and suppress unnecessary reflections. After the first answer is generated (at step $i^{\prime}$), if the detected confidence score $\hat{c}_{i}$ exceeds the early stopping threshold $t_{e}$, the model is explicitly trained to terminate reasoning at that point.
Through this process, the model learns to regulate its generation trajectory based on confidence, strengthening certainty during reasoning and halting once sufficient confidence is reached. As a result, the generation distribution favors precise and concise reasoning chains without compromising task performance.
4 Experiments
Model Method Math-500 GSM8K AIME24 GPQA Average
[…]
DS-7B Origin 90.8 3854.2 100% 93.1 1441.9 100% 54.2 13573.5 100% 51.0 8141.8 100% 72.3 100%
OverThinkSFT 92.2 2537.7 65.8% 93.0 1002.1 69.5% 52.5 11225.3 82.7% 51.1 7638.7 93.8% 72.2 78.0%
SpiritSFT 91.0 2934.6 76.1% 91.4 1107.4 76.8% 51.7 11528.7 84.9% 50.3 7084.3 87.0% 71.1 81.2%
ConCISESFT 92.0 2244.3 58.2% 92.9 831.9 57.7% 52.1 9750.8 71.8% 50.0 5891.5 72.4% 71.8 65.0%
OverThinkSimPO 91.4 2405.1 62.4% 92.9 879.4 61.0% 50.0 9603.2 70.8% 49.9 6304.9 77.4% 71.0 67.9%
SpiritSimPO 87.2 1765.1 45.8% 90.8 687.6 47.7% 38.3 6926.0 51.0% 50.2 5832.0 71.6% 66.6 54.0%
ConCISESimPO 91.0 1945.7 50.5% 92.1 715.1 49.6% 48.3 7744.7 57.1% 48.0 4859.4 59.6% 70.0 54.2%
DS-1.5B Origin 82.2 4783.7 100% 85.4 2219.0 100% 29.2 17464.9 100% 34.5 9492.3 100% 57.8 100%
OverThinkSFT 84.0 3295.6 68.9% 85.2 1199.7 54.1% 30.0 12893.3 73.8% 34.8 8600.0 90.6% 58.5 71.8%
SpiritSFT 83.8 3856.7 80.6% 84.8 1475.9 66.5% 28.8 13684.5 78.4% 32.3 8554.7 90.1% 57.4 78.9%
ConCISESFT 83.6 2700.8 56.5% 84.9 923.3 41.6% 30.0 11359.0 65.0% 35.0 7252.6 76.4% 58.4 59.9%
OverThinkSimPO 83.6 2737.8 57.2% 84.8 1002.9 45.2% 31.26 11465.1 65.7% 34.7 8249.8 86.9% 58.6 63.8%
SpiritSimPO 82.6 2455.3 51.3% 82.6 804.2 36.2% 26.66 9946.3 57.0% 33.1 6910.3 72.8% 56.2 54.3%
ConCISESimPO 83.6 2428.6 50.8% 84.3 803.0 36.2% 30.43 8810.1 50.4% 35.4 7056.1 74.3% 58.4 52.9%
Table 1: Evaluation results of OverThink, SPIRIT, and ConCISE fine-tuned via SFT or SimPO on DeepSeek-R1-Distill 7B and 1.5B, showing accuracy(Acc.), token count(Tok.), and compression ratio(CR) across four benchmarks.
Figure 4: Comparison of StepNum and StepLen across two models and four benchmarks. Stars indicate ConCISE-based methods, which achieve stronger stepwise compression than baselines while maintaining comparable accuracy.
4.1 Settings
Models. We evaluate ConCISE on two Large Reasoning Models (LRMs): DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-1.5B Guo et al. (2025). Although both models exhibit strong mathematical reasoning abilities, their substantial redundancy in reflection steps highlights the need for reasoning chain compression.
Question Set. We find that effective reasoning-chain compression can be achieved with relatively limited data. Based on this observation, we select 2,000 questions from the MATH training setHendrycks et al. (2021) for fine-tuning, ensuring that each question yields a correct answer under greedy decoding and that the set maintains diversity in difficulty levels and response lengths. Detailed analysis is provided in the AppendixB.1.
Evaluation Benchmarks. We evaluate model performance across four benchmarks: GSM8kCobbe et al. (2021), Math-500Hendrycks et al. (2021), AIME24Li et al. (2024), and GPQA_diamondRein et al. (2024). All experiments use the same decoding configuration with temperature = 0.6 and top_p = 0.95Guo et al. (2025). The maximum length is set to 16k for GSM8k, Math-500, GPQA_diamond, and 32k for AIME24. For AIME24 and GPQA_diamond, due to their higher difficulty and smaller sizes, we perform 8 sampling rounds per instance and report the mean values across all metrics to ensure stability and reliability.
Evaluation Metrics. We adopt three primary metrics to comprehensively assess both reasoning accuracy and compression effectiveness. Acc. denotes the final answer accuracy. Tok. refers to the average response length, measured in tokens. CR (Compression Rate) is defined as the ratio of the average response length to that of the original model, with lower values indicating higher compression.
Baseline Methods. We compare our method against existing approaches that aim to remove redundant reasoning and construct efficient reasoning datasets. OverThink.Chen et al. (2024) Samples eight responses from the LRM with a relatively high temperature and selects the shortest correct one. Only the first reflection step after the answer is retained. Spirit.Cui et al. (2025) Iteratively removes steps with minimal impact on perplexity (PPL), deleting 30% of the least critical steps. Adjacent contexts are merged when needed to maintain coherence.
Training Details. We construct training datasets based on the question set using ConCISE, OverThink, and Spirit. For Spirit, compression is applied to reasoning chains generated via greedy search. During training, we adopt two strategies: SFT and SimPO. SFT directly uses the question and its corresponding compressed reasoning trace for supervised fine-tuning. For SimPO, reject samples are also required Rafailov et al. (2023). Following the original OverThink setup, we use the longest correct reasoning chain among the eight samples as the reject samples. To ensure a fair comparison, ConCISE adopts the same strategy. For Spirit, the full reasoning chain before compression serves as reject samples. All models are trained with $lr=1e-6$; SFT runs for 2 epochs and SimPO for 1 epoch.
4.2 Reults and Analysis
This section evaluates multiple compression methods across various reasoning benchmarks. Our results show that ConCISE strikes a superior balance between compression and task performance compared to baseline methods, enabling models to effectively eliminate redundant reasoning steps. Additionally, models fine-tuned with ConCISE on mathematical datasets generalize well to GPQA_diamond, showcasing its robustness.
4.2.1 ConCISE achieves superior compression-performance balance
As shown in Table 1, ConCISE demonstrates strong reasoning chain compression under both SFT and SimPO settings while maintaining comparable accuracy. Under SFT, it achieves a compression rate of ~60%, and under SimPO, it further improves compression to ~50%.
In contrast, OverThink achieves competitive task performance but exhibits weaker compression under SFT. Although it achieves further compression under SimPO, its final compression rate and task performance are comparable to, or even slightly worse than, ConCISE under SFT. As for Spirit, it underperforms in both settings, achieving the weakest compression under SFT and suffering noticeable task performance degradation compared to the other two methods. While Spirit matches ConCISE’s compression rate under SimPO, its accuracy drops more severely, especially on challenging datasets like AIME24. Overall, ConCISE offers a better balance between compression and performance compared to baseline methods.
(a) Step-level analysis of reasoning chains on Math-500
(b) Comparison of training datasets of different methods
Figure 5: Extended analysis of reasoning chain characteristics. More details can be found in the Appendix B.3.
4.2.2 How does ConCISE achieve better compression-performance balance?
To deeply understand how ConCISE achieves a better balance between compression and performance, we perform a stepwise analysis of generated reasoning chains on benchmarks. As shown in Figure 4, we evaluate the average number of reasoning steps in the thinking stage(StepNum) and the average token count per step in the thinking stage(StepLen) across different benchmarks. ConCISE consistently achieves the fewest reasoning steps, while StepLen remains similar to that of the original model. Further details can be found in the AppendixA.3.
Further analysis, focusing on reflection steps in the generated reasoning chains on MATH-500, reveals how each method handles reflection behavior. We use several metrics, including StepNum, Non-RefNum, RefNum, Pre-FAS RefNum, and Post-FAS RefNum. Non-RefNum represents the number of non-reflection steps, and the last three metrics measure the number of reflection steps across the entire reasoning chain, before FAS(First Answer Step), and after FAS, respectively. The results are shown in Figure 5(a).
We find that, in terms of RefNum, Spirit-SFT retains the most reflection steps, followed by OverThink, while ConCISE and Spirit-SimPO retain fewer reflection steps. For Non-RefNum, OverThink and ConCISE perform similarly, while Spirit-SimPO slightly underperforms and also causes a performance drop. This suggests that none of the compression methods can effectively reduce non-reflection steps without compromising model performance. Consequently, it indicates that existing compression methods mainly achieve compression by pruning reflection steps.
For Pre-FAS RefNum, OverThink scores the highest, followed by Spirit, as neither method effectively controls early-stage reasoning. Spirit-SFT, on the other hand, has the highest Post-FAS RefNum, indicating its limited ability to identify and prune redundant steps. In contrast, ConCISE maintains consistent control over both pre- and post-answer reflections. These differences help explain the performance gaps: OverThink’s inability to prune pre-answer reflections limits its compression potential, while Spirit’s failure to identify redundant steps leads to suboptimal pruning, removing essential content and degrading performance. ConCISE, through confidence injection and early stopping, precisely removes redundant reflection steps without harming essential reasoning process.
4.2.3 How does the training dataset affect compression?
Since all methods share identical training configurations and hyperparameters, the quality of the training dataset directly determines the final compression performance. To further investigate which features of the training dataset are highly correlated with the final compression performance, we analyzed the reasoning datasets constructed by ConCISE, OverThink, Spirit, Greedy-Search (reject samples of Spirit), and Longest-of-N (reject samples of ConCISE and OverThink), with the evaluation metrics including Tok., StepNum, StepLen, RefNum, Pre-FAS RefNum, and Post-FAS RefNum. Figure 5(b) shows the distribution of these metrics across the training datasets.
Our results indicate that surface-level metrics such as Tok., StepNum, and StepLen do not directly correlate with the compression rate. For instance, ConCISE and OverThink have nearly identical values on these metrics, but ConCISE achieves much better compression after fine-tuning.
In contrast, other metrics reveal clear distinctions: OverThink records the highest Pre-FAS RefNum, while Spirit leads in Post-FAS RefNum. ConCISE, by comparison, consistently maintains lower values on both. This aligns with the reasoning chain analysis in the previous subsection, indicating that key features from the training data are effectively transferred to the fine-tuned model. These features enable the model to compress reasoning, particularly reflection steps, in line with the dataset design principles.
4.2.4 Can reasoning compression methods generalize across tasks?
Another important question is whether compression methods can generalize beyond the training data distribution. This property is crucial for the practical use of reasoning chain compression.Qu et al. (2025) In our setup, the training dataset is constructed from the question set of the MATH training set. Accordingly, MATH-500 serves as the in-distribution evaluation, GSM8K and AIME24 represent in-domain but out-of-distribution tasks, and GPQA_diamond serves as an out-of-domain task, differing substantially in content, format, and prompt structure.
As shown in Table 1, ConCISE consistently achieves a better balance between compression and performance than baseline methods across both in-domain and out-of-domain tasks. These results indicate that fine-tuning on efficient reasoning chains constructed by ConCISE enables models to structurally suppress redundant reflection steps, a property that holds regardless of task type or prompt format, without compromising overall performance.
Benchmark Method Acc. Tok. StepNum StepLen CR
Math-500 Origin 90.8 3854.2 113.7 33.9 100%
ConCISE 92.0 2244.3 63.0 35.6 58.2%
Conf. Inj. 91.8 2619.6 79.5 33.0 68.0%
Early Stop 92.4 2653.8 68.4 38.8 68.9%
GSM8K Origin 93.1 1441.9 41.4 34.8 100%
ConCISE 92.9 831.9 22.5 37.1 57.7%
Conf. Inj. 92.6 933.8 23.2 40.2 64.8%
Early Stop 92.7 1003.0 22.3 45.0 69.6%
AIME24 Origin 54.2 13573.5 463.0 29.3 100%
ConCISE 52.1 9750.8 333.0 29.3 71.8%
Conf. Inj. 51.3 10165.8 372.2 27.3 74.5%
Early Stop 54.2 12204.6 389.7 31.3 89.9%
Table 2: Ablation study of DeepSeek-R1-Distill-Qwen-7B across three mathematical benchmarks.
4.3 Ablation Study
We conduct an ablation study to evaluate the two core components of ConCISE: Confidence Injection and Early Stopping. For each mechanism, we construct a training dataset and fine-tune the base model using the same configurations as in the main experiments. All experiments are performed in the supervised fine-tuning (SFT) setting and evaluated across three mathematical reasoning tasks: Math-500, GSM8K, and AIME24. The results, summarized in Table 2, show that both mechanisms achieve similar reasoning accuracy to full ConCISE but exhibit weaker compression performance. Both methods focus on reducing redundant reflection steps while maintaining the integrity of the reasoning process.
Confidence Injection suppresses unnecessary iterations by reinforcing the model’s confidence, but it has a limited impact on terminating reasoning after the final answer. Early Stopping, on the other hand, halts excessive post-answer reasoning but does not address earlier stages. Interestingly, Early Stopping nearly preserves the original model’s accuracy on benchmarks like AIME24, suggesting that post-FAS reflection steps often contain substantial redundancy and can be safely stopped without harming task performance. In summary, Confidence Injection reduces redundancy across the entire reasoning sequence, while Early Stopping targets post-answer overthinking. Only by combining both can ConCISE achieve comprehensive redundancy elimination, resulting in superior compression without compromising reasoning quality.
5 Conclusion
In this work, we first modeled the generation of reflection steps in Large Reasoning Models (LRMs) from a confidence perspective. This modeling approach enabled us to systematically identify two primary patterns of redundancy: Confidence Deficit, where models unnecessarily undertrust intermediate steps, and Termination Delay, where reflection persists despite reaching a confident answer. To address these issues, we introduced ConCISE, a novel confidence-guided framework designed to enhance reasoning efficiency. ConCISE employs two targeted mechanisms: Confidence Injection to suppress redundant reflections by boosting internal confidence, and Early Stopping to halt generation upon detecting sufficient answer confidence. Our experiments empirically validate that ConCISE significantly compresses response length (~50% under SimPO) while preserving strong accuracy across diverse reasoning benchmarks. Notably, ConCISE achieves superior compression-accuracy trade-offs over existing methods and exhibits robust generalization to out-of-domain tasks. This study offers a principled approach to modeling and mitigating redundancy in LRM reasoning, enabling more resource-efficient and user-friendly AI systems.
Limitations
While ConCISE demonstrates strong performance in compressing reasoning chains, it is subject to certain limitations that merit further exploration:
Further Compression of Step Length.
While ConCISE effectively targets redundant reflection steps, analysis indicates substantial potential remains for intra-step compression based on average step length. Future work could explore integrating techniques like TokenSkip with ConCISE for greater compression.
Confidence Estimation During Reasoning.
Although ConCISE introduces a lightweight confidence detector for the post-answer phase, it currently lacks a direct mechanism to model confidence prior to the first answer, relying instead on reflection occurrences as an indirect proxy.
Model Dependency and Sensitivity.
Validated on DeepSeek models, ConCISE’s effectiveness and the optimal tuning of its components (e.g., confidence phrases, detector, thresholds) may be sensitive to model architecture or scale. Its performance across diverse LRMs requires further study, potentially necessitating model-specific adaptations.
Appendix A Details about ConCISE
A.1 Confidence Injection
The selection of appropriate confidence phrases is critical for the Confidence Injection component of ConCISE. We curated a pool of 20 distinct phrases, presented in Table 3, by observing the model’s native reasoning expressions, employing manual design, and performing experimental refinement (partly illustrated in Figure 3(a)). The purpose of these phrases is to inject confidence into the model, thereby preventing redundant reflection steps during reasoning. To mitigate potential overfitting to any specific phrase, one phrase is randomly selected from the pool for injection at runtime.
Table 3: The Pool of 20 Confidence Phrases used for Confidence Injection.
Index Confidence Phrase
1 Therefore
2 The reasoning holds
3 Previous steps are correct
4 All steps are valid
5 With this established
6 That sounds reasonable
7 Let’s go ahead
8 Alright, let’s carry on
9 Let’s proceed
10 Let’s progress
11 So, putting it all together
12 The logic stands firm
13 The reasoning process is valid
14 Good, let’s keep going
15 Everything seems reasonable so far
16 This part checks out
17 I think that’s solid. So
18 The reasoning holds, let’s keep going
19 Everything checks out, let’s move on
20 All steps are solid, let’s move forward
A.2 Early Stopping
probing prompt and confidence-indicated tokens. The effectiveness of the Early Stopping mechanism relies heavily on the accuracy of its confidence detector. To identify an appropriate probing prompt, we analyzed the reasoning chains generated by the DeepSeek-R1-Distill-Qwen-7B model using Greedy decoding on the training question set. This analysis revealed that the model frequently expresses confidence during the later part of the reasoning chain. Table 4 summarizes the frequencies of the most common self-reported confidence expressions observed within chains.
Table 4: Frequency of Common Confidence Expressions observed in 2000 Reasoning Chains.
Confidence Expression Frequency Count
I’m confident 423
I feel confident 170
I can be confident that 157
I can confidently say 70
This gives me confidence that 68
I’m pretty confident 43
I’m pretty sure 37
Based on the prevalence of phrases starting with "I’m" leading into confidence statements, we selected "So, I’m" as the fixed probing prompt. Subsequently, we identified "confident", "pretty confident", "sure", and "pretty sure" as the primary confidence-indicative tokens.
The confidence score, $\hat{c}_{i}$, for a given state $S_{i}$ is then calculated by summing the probabilities of generating these tokens or sequences immediately following the probing prompt "So, I’m":
[…]
Here, notation like $P(\text{token}\mid S_{i},\text{sequence})$ represents the conditional probability of generating the next token given the preceding state $S_{i}$ and the specific text sequence provided as context. As demonstrated in Figure 3(b) and supported by our experimental results, the confidence detector designed using this probing prompt and set of confidence-indicative tokens proves effective for the early stopping mechanism in ConCISE.
Early stopping threshold. The selection of the early stopping threshold, $t$, was guided by the principle of tailoring the reflection process to the apparent quality of the reasoning chain. Our objective was to configure the threshold such that:
  • Simple problems with rigorous reasoning steps require no post-answer reflection.
  • Problems of moderate difficulty benefit from exactly one round of reflection.
  • Difficult problems or those with less rigorous reasoning undergo two or more rounds of reflection.
To find a threshold value that aligns with this principle, we empirically evaluated $t_{e}\in\{0.4,0.5,0.6,0.7\}$. For each value, we measured the probability distribution of the stopping points across different stages: stopping after the initial answer generation (0 reflections), after the first reflection, after the second reflection, etc. The results of this analysis are illustrated in Figure 6.
Figure 6: Distribution of early stopping points (probability of stopping at Pre-FAS, FAS, Ref1, 2, …, 5) for different threshold values ($t_{e}$).
Observing the distributions in Figure 6, we found that setting the threshold $t_{e}=0.5$ yielded the stopping behavior that best matched our desired criteria across varying problem difficulties. Therefore, we selected $t_{e}=0.5$ as the definitive early stopping threshold for all subsequent experiments conducted in this work.
A.3 Reflection step detection
The detection of reflection steps, as mentioned throughout this paper, plays a crucial role in both the algorithmic flow and experimental evaluation of ConCISE. Specifically, we employ two distinct methods for identifying these steps, selecting the most suitable one depending on the specific application context. These methods and their corresponding usage scenarios are detailed below.
A.3.1 Rule-based detection.
Reflection steps generated by the model often exhibit distinct linguistic characteristics, frequently containing specific keywords (e.g., "wait", "alternatively") that signal potential reasoning shifts or reconsiderations. Consequently, a straightforward and effective strategy for identifying the start of such steps is rule-based detection. This method checks for the presence of predefined "reflection keywords" within the generated text of a given step, $S_{i}$.
[…]
where $k\subseteq s_{i}$ indicates that the keyword $k$ appears as a substring of step $s_{i}$.The set $K$ of reflection keywords used for rule-based detection consists of the following:
  • "wait", "alternatively", "check", "reconsider", "reflect", "rethink", "reconsidering", "reviewing", "reassess", "pause", "second thought", "reevaluate", "verify".
While effective for identifying the onset of a reflection sequence, this method typically cannot recognize subsequent steps within the same reflection process, as these later reflection steps often lack the initial triggering keywords. Despite this limitation, we utilize this rule-based approach within the Confidence Injection. Since Confidence Injection operates during the model’s generation process, identifying and potentially preventing only the first redundant reflection step is often sufficient to avoid generating the entire reflection. This method is particularly advantageous here due to its efficiency (introducing no computational overhead) and ease of implementation.
Similarly, this rule-based detection is used for the Figure 3(a) analysis to determine if an intervention (like inserting a confidence phrase) triggers the start of a new reflection, as only detecting the initial step with this lightweight method is necessary.
A.3.2 LLM-as-a-Judge
When the task requires identifying all reflection steps within an entire reflection behavior—a capability beyond the rule-based method—a more comprehensive approach is needed. For example, implementing the OverThink baseline necessitates isolating and retaining the complete first reflection sequence that occurs after the FAS(First Answer Step). To address this, we employ an LLM-as-a-Judge methodology.
Specifically, we utilize the Qwen-Max model as the judge. We provide it with both the original problem/question and the model’s complete reasoning chain as input. The prompt supplied to Qwen-Max includes precise definitions of the FAS and the characteristics defining a ’reflection behavior’. The model is instructed to return a structured output that annotates the input chain, identifying the index of the FAS and providing the indices for all detected reflection steps, grouped according to the reflection behavior they belong to. We then parse this structured output to finalize the detection and grouping of reflection steps. Illustrative examples of the prompt structure, input format, and expected response can be found in Figure 9, Figure 10, and Figure 11
Our observations indicate that Qwen-Max performs reliably in identifying FAS and grouping reflection steps when the reasoning chains are of moderate length. However, we noted a decline in performance for chains exceeding 5000 tokens. This degradation is likely attributable to the known limitations of current LLMs in processing very long contexts effectively. Consequently, for the implementation of the OverThink baseline comparison, reasoning chains longer than 3000 tokens were manually annotated to ensure accuracy.
Furthermore, the LLM-as-a-Judge approach is the basis for calculating several reflection-based metrics presented in our results, such as RefNum, Non-RefNum, pre-FAS RefNum, and post-FAS RefNum (shown in Figure 5(a) and Figure 5(b)).
Appendix B Details about Experiments
B.1 Question Set
Figure 7: Difficulty diversity of question set.
Figure 8: Length diversity of question set.
While the construction methodology for the question set is detailed in Section 4, here we focus on illustrating its diversity in terms of difficulty and length.Figures 7 and 8 demonstrate the diversity of the Question set. Difficulty distribution is derived from metric Level in the MATH dataset, while length distribution reflects reasoning chain lengths from DeepSeek-R1-Distill-Qwen-7B under greedy decoding.
B.2 Validation Dataset
For the experiments presented in Figure 3, we curated the Verification Dataset. This was necessary because the required analyses (including First Answer Step (FAS) and reflection detection, as shown in Figure 3(b)) rely on annotations from the LLM-as-a-Judge method (§A.3.2), whose reliability decreases on long reasoning chains. To ensure dependable results, the Verification Dataset comprises another 1000 reasoning chains selected from the MATH training set, filtered based on two criteria: yielding a correct answer under greedy decoding and having a reasoning chain length under 5000. This curated set facilitates reliable LLM-based annotation for the validation experiments.
B.3 Evaluation Dataset
The analysis presented in Figure 5 required LLM-as-a-Judge annotations. To ensure reliable results despite this method’s limitations with long contexts, we filtered the MATH500 dataset. We selected only those problems where reasoning chains generated by all twelve specified finetuning configurations (combinations of ConCISE/OverThink/Spirit methods, 1.5B/7B sizes, and SFT/SimPO techniques) were shorter than 5000 tokens. This yielded a common subset of 340 problems, on which our subsequent comparative analysis of the different methods was exclusively performed.
Prompt You are an AI assistant trained to analyze reasoning steps in a response. Your task has two parts: 1. Examine each reasoning step to determine if it’s part of a reflection process. 2. Identify the earliest step where the final answer (as later shown in boxed{}) is first derived, regardless of whether it is formally written or boxed at that moment. ### [Definition of Reflection]: 1. A reflection process is a sequence of one or more reasoning steps that recheck or doubt a previously made conclusion, such as double-checking calculations, using alternative methods. 2. Typical signals include (but are not limited to): ‘Wait’, ‘Alternatively’, ‘Just to double check’, ’But hold on’, etc. These signals usually mean the start of a new reflection process. 3. However, even without such phrases, if the content of a step reflects a verification or reevaluation, it should be marked as a part of a reflection process. ### [Output Format]: 1. Reflection Step: List all reflection processes as groups of steps. - If Step3 and Step4 form a reflection, write as (Step3, Step4) - If Step5, Step6, Step7 form a new reflection process together, list as a separate group: [(Step3, Step4), (Step5, Step6, Step7)] - Avoid putting a lot of steps into one single reflection process. 2. First Answer Step: Write the earliest step where the final answer is first derived(e.g., Step2). ### [Example]: Question: 2 + 3 = ? Response: Step1: I start with 2 + 3.  Step2: That gives me 5.  Step3: Wait, is that right? Let me make sure…  Step4: But hold on, let me double-check. Maybe I should…  Step5: Wait, no, I think…  Step6: Therefore, the result is 5  Step7: Alternatively, if I use…  Step8: So, the answer is 5  Step9: **Final Answer**: the answer is boxed5. Output: Reflection Step: [(Step3), (Step4), (Step5, Step6), (Step7, Step8)] First Correct Step: Step2 Explanations for Reflection Step: - There is ’wait’, a reflection keyword, in Step3, so Step3 is the start of a reflection process. - There is ’But hold on’ in step4, so step4 means the start of a new reflection process. - Then, there is ’wait’ in step5, so step5 means the start of a new reflection process. - There is ’therefore’ in step6, so (step5,step6) is in the same refleciton process. - There is ’alternatively’ in step7, so step7 is the start of a new reflection process. - There is ’so’ in step8, so (step7,step8) is in the same reflection process. ### [Task]: Now analyze the following question and response: f"Question: {question}" f"Response: {response}" Please output the reflection steps and the first correct step in the format without extra explanation.
Figure 9: Prompt of using the LLM-as-a-Judge method for reflection step detection.
Input Question: What day of the week will it be 40 days from Tuesday? Response: Step1: Alright, so I need to figure out what day of the week it will be 40 days from Tuesday. Hmm, okay. Let me think about how to approach this. Step2: First, I remember that each week has 7 days. So, if something happens every 7 days, it’s on the same day of the week next time. That might be useful here. Maybe I can figure out how many weeks are in 40 days and then see how many extra days that is. Step3: Let me do the division: 40 divided by 7. I think that’s 5 weeks and 5 days. Step4: Wait, let me check. 7 times 5 is 35, right? So, 40 minus 35 is 5. So, 40 days is equal to 5 weeks and 5 days. That makes sense. Step5: So, if today is a Tuesday, then 35 days later is also a Tuesday. Because 35 is exactly 5 weeks, which is 5 times 7. So, after 35 days, it’s still Tuesday. Now, I need to find out what day it is 5 days after Tuesday. Step6: Let me count: Tuesday plus one day is Wednesday, plus two days is Thursday, plus three days is Friday, plus four days is Saturday, and plus five days is Sunday. So, 5 days after Tuesday is $𝑆𝑢𝑛𝑑𝑎𝑦$. Step7: Therefore, 40 days from Tuesday is Sunday. Let me double-check that to make sure I didn’t make a mistake. If today is Tuesday and 7 days later is also Tuesday, then adding another 7 days each time cycles through the week. So, 35 days later, as we calculated, it is Tuesday again. Then, 5 days beyond that would be Sunday. Yeah, that seems right. Step8: Final Answer: The day of the week 40 days from Tuesday is $𝑆𝑢𝑛𝑑𝑎𝑦$.
Figure 10: An example input of using the LLM-as-a-Judge method for reflection step detection.
Output Output: Reflection Step: [(Step4), (Step7)] First Answer Step: Step6
Figure 11: An example output of using the LLM-as-a-Judge method for reflection step detection.

### v2

@@
 Abstract
-Large Reasoning Models (LRMs) perform strongly in complex reasoning tasks via Chain-of-Thought (CoT) prompting, but often suffer from verbose outputs caused by redundant content, increasing computational overhead, and degrading user experience.
-Existing compression methods either operate post-hoc pruning, risking disruption to reasoning coherence, or rely on sampling-based selection, which fails to intervene effectively during generation.
-In this work, we introduce a confidence-guided perspective to explain the emergence of redundant reflection in LRMs, identifying two key patterns: Confidence Deficit, where the model reconsiders correct steps due to low internal confidence, and Termination Delay, where reasoning continues even after reaching a confident answer.
-Based on this analysis, we propose ConCISE (Confidence-guided Compression In Step-by-step Efficient Reasoning), a framework that simplifies reasoning chains by reinforcing the model’s confidence during inference, thus preventing the generation of redundant reflection steps.
-It integrates Confidence Injection to stabilize intermediate steps and Early Stopping to terminate reasoning when confidence is sufficient.
-Extensive experiments demonstrate that fine-tuning LRMs on ConCISE-generated data yields significantly shorter outputs, reducing length by up to ~50% under SimPO, while maintaining high task accuracy.
-ConCISE consistently outperforms existing baselines across multiple reasoning benchmarks.
+Large Reasoning Models (LRMs) perform strongly in complex reasoning tasks via Chain-of-Thought (CoT) prompting, but often suffer from verbose outputs, increasing computational overhead.
+Existing fine-tuning-based compression methods either operate post-hoc pruning, risking disruption to reasoning coherence, or rely on sampling-based selection, which fails to remove redundant content thoroughly.
+To address these limitations, this work begins by framing two key patterns of redundant reflection in LRMs—Confidence Deficit, wherein the model reflects on correct intermediate steps, and Termination Delay, where reflection continues after a verified, confident answer—through a confidence-guided perspective.
+Based on this, we introduce ConCISE (Confidence-guided Compression In Step-by-step Efficient Reasoning), a framework designed to generate concise reasoning chains, integrating Confidence Injection to boost reasoning confidence, and Early Stopping to terminate reasoning when confidence is sufficient.
+Extensive experiments demonstrate that compared to baseline methods, fine-tuning LRMs on ConCISE-generated data yields a better balance between compression and task performance, reducing length by up to ~50% under SimPO, while maintaining high task accuracy.
 ConCISE: Confidence-guided Compression in Step-by-step Efficient Reasoning
-Ziqing Qiao1, Yongheng Deng1, Jiali Zeng2, Dong Wang1, Lai Wei1, Fandong Meng2, Jie Zhou2, Ju Ren1, Yaoxue Zhang1 1Tsinghua University 2Pattern Recognition Center, WeChat AI, Tencent Inc., China
+Ziqing Qiao1thanks: Work was done when Ziqing Qiao was interning at Pattern Recognition Center, WeChat AI, Tencent Inc., Yongheng Deng1thanks: Corresponding authors, Jiali Zeng2, Dong Wang1, Lai Wei1, Guanbo Wang1, Fandong Meng [^2] footnotemark: 2, Jie Zhou2, Ju Ren1, Yaoxue Zhang1 1Department of Computer Science and Technology, Tsinghua University 2Pattern Recognition Center, WeChat AI, Tencent Inc., China qzq24@mails.tsinghua.edu.cn dengyh1013@gmail.comfandongmeng@tencent.com
 1 Introduction
-In recent years, Large Language Models (LLMs) have achieved remarkable progress in natural language understanding and generation, particularly in complex reasoning tasks.
+In recent years, Large Language Models (LLMs) have achieved remarkable progress in natural language processing, particularly in complex reasoning tasks.
 Large Reasoning Models (LRMs), such as OpenAI-o1 Jaech et al.
 (2024) and DeepSeek-R1 Guo et al.
-(2025), which leverage the Chain-of-Thought (CoT) paradigm Wei et al.
-(2022), have demonstrated state-of-the-art performance on benchmarks involving mathematical problem solving and logical inference Muennighoff et al.
+(2025), which leverage the Chain-of-Thought paradigm Wei et al.
+(2022), have demonstrated state-of-the-art performance on benchmarks involving mathematical problem solving and logical reasoningMuennighoff et al.
 (2025); Qwen Team (2024).
@@
 (2025); Chen et al.
+(2024), incuring considerable computational overhead, and posing challenges for deployment in resource-constrained settings Team et al.
 (2025).
-This inefficiency impairs user experience and incurs considerable computational overhead, posing challenges for real-time applications and deployment in resource-constrained settings Team et al.
-(2025).
-Figure 1: Our proposed method, ConCISE, achieves a more favorable trade-off between reasoning compression and task performance than existing methods.
-To mitigate the verbosity of reasoning outputs, recent research has focused on compressing the reasoning chains generated by LRMs Qu et al.
+Figure 1: Training dataset construction workflows: ConCISE (our proposed method) vs. existing methods.
+Figure 2: ConCISE achieves a better trade-off between compression and task performance than baselines.
+To mitigate LRM output verbosity, recent research focuses on compressing their reasoning chains Qu et al.
 (2025); Sui et al.
 (2025).
-A mainstream and effective strategy is finetuning LRMs on concise reasoning datasets, enabling the LRMs to generate shorter reasoning chains directly during inference Ma et al.
+A prominent strategy involves fine-tuning LRMs on concise reasoning datasets, enabling them to generate shorter responses Ma et al.
 (2025); Chen et al.
@@
 (2025).
-The quality of such compression heavily relies on how these datasets are constructed.
-One common approach is sampling-based selection, which generates multiple candidate reasoning chains and selects the shortest correct one Team et al.
-(2025), or removes redundant reflection steps after the final answer is obtained Chen et al.
-(2024).
-While simple, these approaches lack control over generation and rely on the reasoning chains naturally generated by the model, which may still include unnecessary steps, reducing compression effectiveness.
-Another line of work relies on post-hoc pruning, where redundant or less informative reasoning steps are identified and removed from a complete reasoning chain Cui et al.
+The effectiveness of this compression strategy hinges on the training dataset design.
+As shown in Figure 1, common sampling-based selection, which generates multiple candidates and picks the shortest correct one, or further removes post-answer redundant reflections Team et al.
+(2025); Chen et al.
+(2024), lacks control during generation, potentially leaving unnecessary steps and reducing compression effectiveness.
+Another approach, post-hoc pruning, identifies and removes redundant or less important steps from reasoning chains Cui et al.
 (2025); Xia et al.
-(2025).
-However, such modifications can disrupt the continuity of reasoning, leading to degraded performance of LRMs after fine-tuning.
-To address the limitations of existing approaches, we seek to construct reasoning chains that are both compact and logically intact, effectively removing redundant reflection steps without sacrificing the performance of LRMs after fine-tuning.
-To this end, we analyse the reasoning behavior of LRMs and identify that the reflection behavior is not solely determined by correctness, but is heavily influenced by the model’s internal confidence.
-In particular, insufficient confidence can cause the model to re-evaluate correct steps, resulting in redundant reflections.
-We thus adopt a confidence-guided perspective to explain why LRMs generate reflection steps.
-From this view, we identify two key patterns underlying redundant reflection: Confidence Deficit, where the model reconsiders correct intermediate steps due to low internal confidence; and Termination Delay, where reflection continues even after a confident final answer is reached.
-These patterns substantially inflate reasoning chains and offer actionable insights for constructing concise reasoning data.
-Building on this analysis, we propose ConCISE (Confidence-guided Compression In Step-by-step Efficient reasoning)—the first framework that leverages a confidence-guided perspective to actively construct concise reasoning datasets by suppressing redundant reflection during generation.
-ConCISE consists of two complementary components—Confidence Injection and Early Stopping—designed to mitigate Confidence Deficit and Termination Delay, respectively.
-Confidence Injection inserts confidence phrases before potential reflection points to strengthen the model’s belief in intermediate steps and suppress unnecessary reflection.
-Early Stopping employs a lightweight confidence detector to monitor the model’s internal confidence and terminate reasoning once sufficient confidence is achieved.
-Through the synergy of these two mechanisms, ConCISE enables a more efficient and compact reasoning chain.
-We then fine-tune LRMs using ConCISE-generated data via SFT Wei et al.
+(2025), risks disrupting reasoning coherence and degrading the performance of LRMs after fine-tuning.
+To overcome existing limitations, we aim to construct compact, coherent reasoning chains as training datasets by precisely removing redundant reflections, ensuring LRMs do not suffer performance degradation after fine-tuning.
+To this end, based on the understanding that reflections are not solely determined by correctness Yang et al.
+(2025), we thus adopt a confidence-guided perspective to understand the generation of reflection steps in LRM’s reasoning processes.
+This perspective offers a clear interpretation of two key patterns of redundancy: Confidence Deficit, where low internal confidence causes models to undertrust and reflect on correct intermediate steps; and Termination Delay, where reflection persists despite a repeatedly verified answer.
+These patterns inflate reasoning chains and provide actionable insights for how to create concise reasoning chains.
+Therefore, we propose ConCISE (Confidence-guided Compression In Step-by-step Efficient reasoning), a framework that leverages a confidence-guided perspective for constructing concise reasoning data by actively suppressing redundant reflection during generation.
+ConCISE features two complementary components designed to mitigate Confidence Deficit and Termination Delay, respectively: Confidence Injection, which inserts tailored phrases before potential reflection points to boost the model’s internal confidence, consequently curbing unnecessary reflection; and Early Stopping, which employs a lightweight confidence detector to monitor internal confidence after an answer is reached, terminating reasoning upon achieving sufficient confidence.
+The synergy of these mechanisms enables ConCISE to produce more efficient and compact reasoning chains.
+We then fine-tune four mainstream LRMs using ConCISE-generated data via SFT Wei et al.
 (2021) and SimPO Meng et al.
-(2024), and evaluate their performance on multiple reasoning benchmarks.
-Experimental results show that ConCISE achieves a superior trade-off between reasoning compression and task performance.
-As shown in Figure 1, ConCISE reduces average response length by ~50% under SimPO while maintaining high accuracy; under SFT, it also significantly outperforms existing methods in compression efficiency.
-These results demonstrate that LRMs fine-tuned with ConCISE-generated data learn to avoid generating redundant reflection steps, achieving efficient reasoning without sacrificing performance.
-The main contributions of this work are summarized as follows:
-  • We present the first confidence-guided explanation of redundant reflection in LRM reasoning, identifying key patterns that underlie reasoning verbosity.
-  • We propose ConCISE, a framework that proactively suppresses redundant reflection during inference and constructs high-quality, concise reasoning data for LRM fine-tuning.
-  • Through extensive evaluation on multiple reasoning benchmarks, we demonstrate that LRMs fine-tuned with ConCISE-generated data achieve a better trade-off between reasoning chain compression and task performance.
+(2024), evaluating their performance across multiple reasoning benchmarks.
+As shown in Figure 2, Experimental results demonstrate that ConCISE consistently achieves a superior trade-off between reasoning compression and task performance under both SFT and SimPO settings compared to baselines.
+Notably, ConCISE reduces average response length by approximately 50% under SimPO while maintaining high accuracy.
+Our subsequent analysis reveals that ConCISE’s unique training data design enables LRMs, after fine-tuning, to learn to strategically avoid generating redundant reflection steps without harming essential critical reasoning content, thereby achieving efficient reasoning compression while task performance is well maintained.
 2 Related Work
-Figure 2: The ConCISE reasoning pipeline applied to solving the equation $73x-2=$ 7.
-The figure compares the original 8-step reasoning process with the streamlined 5-step process generated by ConCISE.
-It illustrates how ConCISE models reflection based on confidence values, identifying Confidence Deficit and Termination Delay, and applying Confidence Injection and Early Stopping to suppress redundant reasoning.
-Here, $C_{i}$ denotes the internal confidence of each step, and $T_{i}$ represents its corresponding threshold.
-Recent research has increasingly focused on mitigating verbosity and redundancy in reasoning chains generated by Large Reasoning Models (LRMs)Qu et al.
-(2025).
-While CoT enables step-by-step reasoning, it often produces long outputs that increase latency, computational costs, and may degrade final accuracyWu et al.
+Figure 3: Illustration of ConCISE’s confidence-guided approach: identifying patterns (Confidence Deficit, Termination Delay) and applying mechanisms (Confidence Injection, Early Stopping) to suppress redundant reflections, shown in contrast to the original reasoning process.
+$C_{i}$ denotes step confidence and $T_{i}$ its threshold.
+Recent research has increasingly focused on mitigating verbosity and redundancy in reasoning chains generated by LRMs, which often produces long outputs that increase computational costs, and even degrade accuracy Wu et al.
 (2025); Nayab et al.
@@
 (2025).
-To address this, existing approaches can be broadly categorized into three types: input-based methods, decoding-time interventions, and model-based training strategies Sui et al.
+To address this problem, existing approaches can be broadly categorized into three classes: Input-based, Decoding-based, and Model-based Sui et al.
 (2025).
-Input-based Methods aim to promote concise reasoning by modifying input prompts without changing the model architecture or decoding strategyLee et al.
+Input-based Methods aim to promote concise reasoning by modifying input texts Lee et al.
 (2025).
-Common techniques include imposing token limits or instructing the model to reason briefly Han et al.
+Common techniques include imposing token limits in prompts or instructing the model to reason briefly Han et al.
 (2024); Renze and Guven (2024).
-However, these approaches often fail when the model does not consistently follow the instructions.
+However, these approaches often fail when the model does not consistently follow these instructions.
 Another approach, task routing, dynamically selects an appropriate model or reasoning strategy based on input characteristics like question complexity Ong et al.
@@
 Its effectiveness, however, depends on the accurate assessment of input characteristics, which is not always reliable.
-Decoding-based Methods intervene during inference to control the reasoning process.
-One direction compresses intermediate steps into latent representations rather than explicit text, which improves brevity but sacrifices interpretability Hao et al.
+Decoding-based Methods intervene during decoding to control the reasoning process.
+One common approach compresses steps into latent representations rather than explicit text, which improves brevity but sacrifices interpretability Hao et al.
 (2024); Shen et al.
 (2025b).
-Another strategy uses dynamic decoding, where lightweight models or heuristics evaluate each reasoning step and decide whether to retain, modify, or discard it during generation Sun et al.
+Another strategy uses dynamic decoding, evaluating each reasoning step and deciding whether to modify or discard it during generation Sun et al.
 (2024); Zhang et al.
@@
 (2025).
-While effective at reducing verbosity, these methods introduce additional computational overhead and may disrupt the reasoning flow.
+While effective at reducing verbosity, these methods introduce additional computational overhead and may disrupt the reasoning coherence.
 Model-based Methods train models to generate concise reasoning directly.
@@
 (2025a); Aggarwal and Welleck (2025); Luo et al.
-(2025); Arora and Zanette (2025), but such methods can be unstable, sensitive to reward formulation, and impose significant training costs.
-More commonly, LRMs are fine-tuned on datasets with concise reasoning chainsMa et al.
+(2025); Arora and Zanette (2025); Yu et al.
+(2025), but can be sensitive to reward formulation and impose significant costs.
+Another common approach is to fine-tune LRMs on datasets with concise reasoning chainsMa et al.
 (2025); Cui et al.
-(2025); Liu et al.
 (2025); Xia et al.
 (2025).
-These datasets are often created by selecting the shortest correct chains or removing redundant parts post-hoc.
-Although effective, these methods lack a principled understanding of redundancy, potentially retaining unnecessary steps or removing useful context, which may hurt compression efficiency or model performance.
-Motivated by the limitations of existing methods, we propose a model-based framework ConCISE.
-Specifically, we achieve precise identification and suppression of redundant reflection steps throughout the entire reasoning process by analyzing and explaining reflection behavior from a confidence-guided perspective.
-This mechanism allows the model to generate concise and efficient reasoning chains while maintaining the flow of reasoning.
-By doing so, we construct a high-quality training dataset and effectively transfer this capability to the model through fine-tuning, enabling it to perform concise reasoning without compromising model performance.
+These datasets are often created by selecting the shortest correct chain or removing redundant parts post-hoc.
+Such methods may inadvertently retain redundant reflection steps or remove useful context, degrading compression efficiency or model performance.
+Motivated by the limitations of existing methods, we propose ConCISE, a model-based method that precisely identifies and actively suppresses redundant reflection steps throughout the reasoning process.
+This approach enables the model to generate efficient reasoning chains while preserving reasoning coherence.
+Leveraging ConCISE, we construct a high-quality training dataset and fine-tune the model to adopt this concise reasoning ability without compromising model performance.
 3 Method
-3.1 Confidence-guided Formulation of Reflection in LRM Reasoning
-We observe that reflection in LRMs is not solely triggered by correctness judgments.
-In many cases, reflection steps are invoked even on correct reasoning steps, including those that have already been verified, leading to substantial redundancy.
-This suggests that reflection behavior is also linked to the model’s internal confidence rather than correctness alone.
-To explain this phenomenon, we adopt a confidence-guided perspective to formalize when and why LRMs engage in reflection.
-Let $S_{i}={s_{1},s_{2},\ldots,s_{i}}$ denote the partial reasoning chain up to step $i$, where each $s_{j}$ is a textual reasoning unit.
-We associate each step $s_{i}$ with a confidence score $c_{i}\in[0,1]$, representing the model’s internal belief in the correctness of that step.
-The model’s generation policy $\pi_{\theta}$ maps the current reasoning context $S_{i}$ to the next step $s_{i+1}$.
-To model the decision between proceeding and reflecting, we introduce a dynamic threshold $t_{i}\in[0,1]$, which may vary with the model architecture or the evolving context.
-At each step, the model compares its internal confidence $c_{i}$ with this threshold $t_{i}$ to determine whether to continue reasoning or enter reflection.
-Formally:
-$s_{i+1}=\pi_{\theta}(S_{i})=\text{ReflectionStep}\quad\text{if }c_{i}<t_{i}$
-Guided by this formulation, we analyze the reflective behavior of LRMs and formally characterize two recurrent patterns responsible for reflection-related redundancy: Confidence Deficit and Termination Delay.
-These patterns explain how the model’s internal confidence leads to the occurrence of redundant reflection.
-We will elaborate on each in the following subsection.
-(a) Reflection probability after confidence injection.
-(b) Detected confidence progression during reasoning.
-Figure 3: Validation experiments for Confidence Injection and Termination Delay. The dataset used in these experiments is described in Appendix B.2.
+3.1 Confidence-guided Formulation
+Reflections in LRMs are not solely triggered by correctness; in many cases, reflection steps are invoked even on correct steps that have been verified Yang et al.
+(2025).
+This suggests that reflection behavior is also linked to the model’s internal confidence about current reasoning rather than correctness alone.
+To explain this, we adopt a confidence-guided perspective to formalize when and why LRMs engage in reflection.
+Let $S_{i}=\{s_{1},s_{2},\ldots,s_{i}\}$ denote the partial reasoning chain up to step $i$, where each $s_{i}$ is a textual reasoning unit.
+We associate each step $s_{i}$ with a confidence score $c_{i}\in[0,1]$, representing the model’s internal confidence of that step.
+The generation policy of the LRM, denoted by $\pi_{\theta}$, maps the current reasoning context $S_{i}$ to the next step $s_{i+1}$.
+To model the decision between proceeding and reflecting, we introduce a dynamic threshold $t_{i}\in[0,1]$, which may vary with the model or the context.
+At each step, the model generates a ReflectionStep as $s_{i+1}$ if its current internal confidence $c_{i}$ falls below the threshold $t_{i}$ (i.e., $c_{i}<t_{i}$).
+From this perspective, we further analyze the reflection behavior of LRMs and formally articulate two key patterns responsible for reflection-related redundancy existing in the reasoning process: Confidence Deficit and Termination Delay.
 3.2 Two Key Patterns of Redundancy
@@
 One major source of redundancy in LRMs stems from their tendency to undertrust their correct intermediate steps.
-LRMs often display unexpected hesitation despite exhibiting fine-grained reasoning capabilities and achieving higher stepwise accuracy, triggering reflection even on simple and unambiguous reasoning steps.
-This self-undermining behavior leads to redundant reflection with minimal semantic gain, ultimately increasing computational cost and disrupting the efficiency of the reasoning process.
+LRMs often display unexpected reflection despite exhibiting fine-grained reasoning capabilities and achieving high stepwise accuracy, triggering reflection even on simple and unambiguous reasoning steps.
+This self-undermining behavior leads to redundant reflection with minimal semantic gain.
 We refer to this phenomenon as Confidence Deficit.
@@
 Termination Delay.
-Beyond Confidence Deficit, LRMs exhibit another redundant reasoning pattern we term Termination Delay: After producing a confident final answer, the model is expected to conclude with minimal additional reasoning.
+LRMs exhibit another important redundant reasoning pattern we term Termination Delay: After producing a confident final answer, the model is expected to conclude with minimal additional reasoning.
 However, it often continues to generate unnecessary reflection steps even after repeatedly verifying the same conclusion.
-This behavior is attributed to a sharp increase in the confidence threshold $t_{i}$ after reaching the answer at step $i^{\prime}$, making it increasingly difficult for the model’s internal confidence $c_{i}$ to exceed $t_{i}$.
-As a result, even when $c_{i}$ becomes relatively high, it may still fall short of the heightened $t_{i}$, preventing the model from terminating, leading to verbose post-answer reasoning.
+This behavior can be attributed to a sharp increase in the confidence threshold $t_{i}$ after reaching the answer at step $i^{\prime}$, making it increasingly difficult for the model’s internal confidence $c_{i}$ to exceed $t_{i}$.
+As a result, even when $c_{i}$ becomes relatively high, it may still fall short of the heightened $t_{i}$, leading to verbose post-answer reasoning.
 Formally, Termination Delay can be characterized as:
@@
 3.3 The ConCISE Framework
-To mitigate the above patterns—Confidence Deficit and Termination Delay—we propose ConCISE, an integrated framework that dynamically steers the reasoning generation process to reduce redundant reflections and produce concise reasoning chains.
-ConCISE integrates two core mechanisms: Confidence Injection, which proactively inserts confidence phrases to suppress unnecessary reflection steps during reasoning, and Early Stopping, which halts further generation once sufficient post-answer confidence is detected.
-Together, these mechanisms operate in synergy to generate concise reasoning chains without compromising inference quality.
+To mitigate the above patterns—Confidence Deficit and Termination Delay—we propose ConCISE, a framework that dynamically steers the reasoning generation process to reduce redundant reflections and produce concise reasoning chains.
+ConCISE integrates two mechanisms: Confidence Injection, which actively inserts confidence phrases to suppress unnecessary reflection steps, and Early Stopping, which halts generation once sufficient post-answer confidence is detected.
+These two mechanisms operate together to generate concise reasoning chains without compromising inference quality.
 The overall process is illustrated in Figure 3.
@@
 The key idea is to boost the model’s internal confidence during reasoning by inserting designed phrases, referred to as confidence phrases, to prevent unnecessary reflection steps.
-Considering that indiscriminate injection could disrupt reasoning, we selectively insert confidence phrases at critical points where the model’s confidence is relatively low—that is, when $c_{i}<t_{i}$.
-Specifically, at each reasoning step $s_{i}$, the model first generates the next step $s_{i+1}=\pi_{\theta}(S_{i})$, where $S_{i}$ denotes the reasoning chain up to step $i$.
-If $s_{i+1}$ is identified as a reflection step, indicating insufficient confidence, we retroactively modify the input by appending a confidence phrase $p_{i}$ sampled from a curated pool $\mathcal{P}$, and regenerate $s_{i+1}$.
+Considering that indiscriminate injection could disrupt reasoning, we selectively insert confidence phrases at critical points where the model’s confidence is relatively low, that is, when $c_{i}<t_{i}$, a condition that signifies an impending reflection step.
+Specifically, at each step $s_{i}$, the model first generates $s_{i+1}=\pi_{\theta}(S_{i})$.
+If $s_{i+1}$ is a reflection step, we retroactively modify the input $S_{i}$ by appending a confidence phrase $p_{i}$ sampled from a curated pool $\mathcal{P}$, and regenerate $s_{i+1}$.
 Formally, the updated generation process is defined as:
-$s_{i+1}=\pi_{\theta}(S_{i})\quad\text{if}\;c_{i}\geq t_{i}\;\text{else}\quad\pi_{\theta}(S_{i},p_{i})$
-We manually constructed the initial version of confidence phrase pool $\mathcal{P}$ based on observations of certainty in original reasoning chains.
-To ensure generalization, the pool includes diverse formulations, and confidence phrases are randomly sampled from it during inference.
-To evaluate and further refine the confidence phrase pool, we conducted experiments measuring the effectiveness of different phrases.
-Specifically, we inserted each candidate phrase at reflection positions and recorded the probability that the model still generated a reflection step.
-This probability indicates whether the updated confidence $c_{i}^{\prime}$ (after phrase insertion) exceeds the dynamic threshold $t_{i}$.
-As shown in Figure 3(a), different phrases exhibit varying levels of effectiveness, some notably reduce the likelihood of reflection, while others have only a marginal impact.
-Notably, even the best-performing phrases still yield a reflection rate above 20%, suggesting that confidence injection reduces redundant reflections while retaining necessary verification.
-Based on these results, we selected the top 20 phrases with the lowest reflection rates to form the final confidence phrase pool.
-The composition of the phrase pool and further experimental details are provided in Appendix A.1.
+$s_{i+1}=\pi_{\theta}(S_{i})\quad\text{if}\;c_{i}\geq t_{i}\;\text{else}\;\pi_{\theta}(S_{i},p_{i})$
+We manually constructed the initial version of $\mathcal{P}$.
+To evaluate and further refine the initial pool, we conducted experiments measuring the effectiveness of different phrases.
+We inserted each candidate phrase before the reflection steps and recorded the probability that the model still generated a reflection step.
+This probability indicates whether the updated confidence $c_{i}^{*}$ (after phrase insertion) exceeds the dynamic threshold $t_{i}$.
+As shown in Figure 4(a), different phrases exhibit varying reflection rates.
+Notably, even the best-performing phrases still yield a reflection rate around 20%, suggesting that confidence injection reduces redundant reflections while still retaining necessary verifications.
+We then selected 20 phrases with the lowest reflection rates to form the final confidence phrase pool.
+The composition of the phrase pool and detection of reflection steps are provided in Appendix A.1.
 3.3.2 Early Stopping
-Although Confidence Injection effectively boosts the model’s confidence during reasoning, it does not fully mitigate the pattern of Termination Delay, which is caused by the significant increase in the dynamic threshold $t_{i}$ after the model generates an answer, leading the model to continue unnecessary reflections even when its confidence is already high.
+(a) Reflection probability of next step after phrase injection.
+(b) Confidence calculated by our detector.
+Figure 4: Effectiveness and necessity of Confidence Injection and Termination Delay(details in Appendix B.1).
+Although Confidence Injection effectively boosts the model’s confidence during reasoning, it does not fully mitigate Termination Delay, which is caused by the high threshold $t_{i}$ after the model reaches an answer, leading the model to continue unnecessary reflections even when its confidence is already high.
 To address this, we design an Early Stopping mechanism based on direct estimation of the model’s internal confidence.
-Specifically, we construct a lightweight confidence detector to provide a quantitative proxy for the model’s internal confidence at each step after the first answer is generated.
-We introduce a probing prompt (such as "So I’m") after the current reasoning context and analyze the probability distribution of generated continuations.
-Based on statistical analysis, we collect a set of confidence-indicative tokens $\mathcal{W}_{+}$, which reflect affirmations of high certainty.
-Formally, the detected confidence score $\hat{c}_{i}$ at step $i$ is calculated as:
+Specifically, we construct a lightweight confidence detector to provide a quantitative proxy for the model’s internal confidence about the answer.
+We introduce a probing prompt after the current reasoning context and analyze the probability distribution of generated continuations.
+Based on the statistical analysis shown in Appendix A.2, we collect a set of confidence-indicative phrases $\mathcal{W}_{+}$, which reflect affirmations of high certainty.
+The detected confidence score $\hat{c}_{i}$ at step $i$ is calculated as:
 $\hat{c}_{i}=\sum_{w\in\mathcal{W}_{+}}p(w\mid S_{i},\text{Probing prompt}),$
-To validate the confidence detector and examine how model confidence evolves post-FAS(First Answer Step), we calculate the average detected confidence $\hat{c}_{i}$ across key reasoning stages, including FAS and the subsequent five reflection steps (Ref1 to Ref5).
-As shown in Figure 3(b), the model exhibits low confidence at the FAS, but after the first reflection, confidence increases sharply and continues to rise through subsequent reflections.
-Notably, the final reflection step has an average detected confidence of $0.93$, suggesting that for $i>i^{\prime}$, the dynamic threshold $t_{i}$ effectively stabilizes around this high value, motivating the introduction of a manually controlled lower threshold.
-Therefore, we define an early stopping threshold $t_{\text{e}}=0.5$.
-The model is instructed to terminate generation as soon as its detected confidence $\hat{c}_{i}$ exceeds $t_{\text{e}}$.
-Further details regarding the selection of $t_{e}$ and the confidence detector are provided in Appendix A.2.
-In addition, we also compute the average and maximum $\hat{c}_{i}$ before the FAS (Avg Pre-FAS and Max Pre-FAS), both of which remain close to zero.
-This indicates that our confidence detector only becomes effective after the first answer is generated, ensuring that premature termination during early reasoning stages does not occur.
+where $p(\cdot|\cdot)$ denotes the continuation probabilities determined by $\pi_{\theta}$.
+We then calculate the average $\hat{c}_{i}$ after the First Answer Step (FAS) is generated, including the subsequent five reflection steps (Ref1-5).
+As shown in Figure 4(b), the model exhibits relatively low confidence at the FAS, but increases sharply after the first reflection, and continues to rise through subsequent reflections.
+Notably, when the reasoning chain stops, the average confidence is $0.95$, motivating the introduction of a manually controlled lower threshold $t_{e}$.
+We ultimately set $t_{e}=0.5$ through experiments.
+Additionally, the application of the threshold is carefully controlled to prevent premature termination: once the detected confidence $\hat{c}_{i}$ exceeds $t_{e}$, the model is first prompted to output a final answer (e.g., by appending Final Answer:).
+The reasoning process halts only if this answer is subsequently verified as correct; otherwise, generation continues.
+More details regarding the selection of $t_{e}$ and Early Stopping are provided in Appendix A.2.
 3.4 Building Efficient Reasoning Chains
+Preparation: LRM generation policy: $\pi_{\theta}$, Confidence Phrase Pool: $\mathcal{P}$, Early Stopping Threshold $t_{e}$, Prompt Template $T$
+Input: Question $q$, Ground Truth gtgt
+Initialize: Reasoning chain $S_{0}\leftarrow T(q)$
+for each reasoning step $i=1,2,\ldots$ do
+ $s_{i}\leftarrow\pi_{\theta}(S_{i-1})$
+ if $s_{i}$ is a reflection step then
+      Sample a confidence phrase $p_{i}\in\mathcal{P}$ $s_{i}\leftarrow\pi_{\theta}(S_{i-1},p_{i})$
+ $S_{i}\leftarrow S_{i-1}+s_{i}$
+   Compute detected confidence $\hat{c}_{i}$
+ if $\hat{c}_{i}>t_{e}$ then
+    $a\leftarrow\pi_{\theta}(S_{i},\text{'Final Answer:'})$
+    if $\text{isequal}(a,gt)$ then
+       break
+if $\text{isequal}(a,gt)$ then
+   Summary: $S\leftarrow S_{i}+\pi_{\theta}(S_{i},^{\prime}\text{</think>}^{\prime})$
+ Output: ConCISE reasoning chain $S$
+else
+ $S_{i}$ and $q$ are discarded
+Algorithm 1 Workflow of ConCISE.
+Model Math-500 GSM8K AIME24 GPQA Average
+[…]
+DeepSeek-7BOrigin 90.8 3854 100% 93.1 1442 100% 54.2 13574 100% 51.0 8142 100% 72.3 100%
+OverThinkSFT 92.2 2538 66% 93.0 1002 70% 52.5 11225 83% 51.1 7639 94% 72.2 78%
+SpiritSFT 91.0 2935 76% 91.4 1107 77% 51.7 11529 85% 50.3 7084 87% 71.1 81%
+ConCISESFT 92.0 2244 58% 92.9 832 58% 52.1 9751 72% 50.0 5892 72% 71.8 65%
+OverThinkSimPO 91.4 2405 62% 92.9 879 61% 50.0 9603 71% 49.9 6305 77% 71.0 68%
+SpiritSimPO 87.2 1765 46% 90.8 688 48% 38.3 6926 51% 50.2 5832 72% 66.6 54%
+ConCISESimPO 91.0 1946 51% 92.1 715 50% 48.3 7745 57% 48.0 4859 60% 70.0 54%
+DeepSeek-1.5BOrigin 82.2 4784 100% 85.4 2219 100% 29.2 17465 100% 34.5 9492 100% 57.8 100%
+OverThinkSFT 84.0 3296 69% 85.2 1200 54% 30.0 12893 74% 34.8 8600 91% 58.5 72%
+SpiritSFT 83.8 3857 81% 84.8 1476 67% 28.8 13685 78% 32.3 8555 90% 57.4 79%
+ConCISESFT 83.6 2701 57% 84.9 923 42% 30.0 11359 65% 35.0 7253 76% 58.4 60%
+OverThinkSimPO 83.6 2738 57% 84.8 1003 45% 31.3 11465 66% 34.7 8250 87% 58.6 64%
+SpiritSimPO 82.6 2455 51% 82.6 804 36% 26.7 9946 57% 33.1 6910 73% 56.2 54%
+ConCISESimPO 83.6 2429 51% 84.3 803 36% 30.4 8810 50% 35.4 7056 74% 58.4 53%
+Skywork-7BOrigin 93.6 4178 100% 93.2 2111 100% 62.9 12464 100% 51.1 8374 100% 75.2 100%
+OverThinkSFT 92.8 3410 82% 93.1 1561 74% 59.2 11047 89% 50.7 8128 97% 73.9 85%
+SpiritSFT 93.4 3279 79% 93.1 1452 69% 58.8 11815 95% 52.0 7565 90% 74.2 83%
+ConCISESFT 93.2 2740 66% 92.7 1247 59% 59.2 9871 79% 51.8 6543 78% 74.2 71%
+OverThinkSimPO 93.8 2758 66% 93.4 1114 53% 56.3 9422 76% 51.5 6773 81% 73.7 69%
+SpiritSimPO 92.0 2329 56% 92.3 801 38% 52.5 8055 65% 51.0 6663 80% 72.0 59%
+ConCISESimPO 93.0 2207 53% 93.3 882 42% 55.9 7598 61% 51.7 5668 68% 74.0 56%
+Qwen3-8BOrigin 93.4 5142 100% 95.6 2211 100% 73.3 15094 100% 60.6 7309 100% 80.7 100%
+OverThinkSFT 93.2 4963 97% 95.5 2133 96% 72.5 14123 94% 60.3 7204 99% 80.4 96%
+SpiritSFT 93.2 4184 81% 95.2 1675 76% 73.3 13854 92% 60.0 6784 93% 80.4 85%
+ConCISESFT 93.0 4712 92% 95.6 1907 86% 72.1 14168 94% 60.4 6952 95% 80.6 92%
+OverThinkSimPO 93.6 2488 48% 95.2 929 42% 69.2 9468 63% 57.1 3524 48% 78.9 50%
+SpiritSimPO 93.2 2920 57% 95.3 997 45% 67.9 10441 69% 56.8 3473 48% 78.3 55%
+ConCISESimPO 93.0 2271 44% 95.5 841 38% 71.6 10098 67% 57.6 3751 51% 79.4 50%
+Table 1: Comparison of OverThink, SPIRIT, and ConCISE methods, fine-tuned with SFT or SimPO on four LRMs.
+Metrics include Accuracy (Acc.), Token Count (Tok.), and Compression Ratio (CR) over four benchmarks.
+Light gray rows denote original LRM performance.
+Best and second-best results are in bold and underlined respectively.
+Note that the “Average” column is computed as an unweighted mean across benchmarks, and this convention applies to all subsequent tables as well.
 ConCISE integrates Confidence Injection and Early Stopping to dynamically control the reasoning generation process.
 Given an input question, the LRM generates the reasoning chain step-by-step.
-At each step, the model first examines whether the next generated step is a reflection step.
-If a reflection is detected, Confidence Injection is applied by appending a confidence prompt and regenerating the step with reinforced internal confidence.
-After generating the next reasoning step, a lightweight confidence detector estimates the detected confidence score $\hat{c}_{i}$.
-If $\hat{c}_{i}$ exceeds a predefined threshold $t_{e}$, the model terminates generation, assuming sufficient confidence in the answer.
-Otherwise, the generated step is appended to the current reasoning chain, and the process repeats..
-Algorithm 1 summarizes the overall coordinated reasoning process in ConCISE.
-Algorithm 1 Algorithmic workflow of ConCISE for generating concise reasoning chains.
-1:Preparation:
-2:    Large Reasoning Model $\pi_{\theta}$
-3:    Confidence Phrase Pool $\mathcal{P}$
-4:    Early Stopping Threshold $t_{e}=0.5$
-5:    Prompt Template $T$
-6:Input: Question $Q$
-7:Initialize: Reasoning chain $S_{0}=T(Q)$
-8:for each reasoning step $i=1,2,\ldots$ do
-9:     Generate $s_{i}=\pi_{\theta}(S_{i-1})$
-10:     if $s_{i}$ is a reflection step then
-11:         Apply Confidence Injection:
-12:             Sample a confidence phrase $p_{i}\in\mathcal{P}$
-13:             Regenerate $s_{i}=\pi_{\theta}(S_{i-1}+p_{i})$
-14:     end if
-15:     Compute detected confidence $\hat{c}_{i}$
-16:     if $\hat{c}_{i}>t_{e}$ then
-17:         Terminate
-18:     end if
-19:     Accept $s_{i}$ and update $S_{i}=S_{i-1}+s_{i}$
-20:end for
+At each step $s_{i}$: if it is identified as a reflection step, Confidence Injection is applied by appending a confidence prompt and regenerating $s_{i}$ with boosted internal confidence.
+$s_{i}$ is then added to the current chain $S_{i}$.
+Subsequently, a lightweight confidence detector estimates the detected confidence score $\hat{c}_{i}$.
+If $\hat{c}_{i}>t_{e}$, the LRM is prompted to provide a final answer $a$.
+Crucially, the generation process terminates early at this stage if, and only if, this answer $a$ matches the ground truth $gt$ and $\hat{c}_{i}>t_{e}$, otherwise reasoning continues to the next step.
+After the loop concludes, should a correct final answer $a$ have been verified, a concluding summary (e.g., prompted by </think>) is generated and appended to the accumulated steps $S_{i}$ to form the complete ConCISE reasoning chain $S$, otherwise, if a wrong answer was finally achieved, the partial chain and the original question are discarded.
+Algorithm 1 summarizes this coordinated reasoning process in ConCISE.
 3.5 Training Objective and Fine-tuning
-To align LRM generation with the behavior encouraged by ConCISE, we first construct a fine-tuning dataset by applying the ConCISE pipeline to a set of reasoning questions.
-This yields concise reasoning traces in which redundant reflections are actively suppressed and timely termination is encouraged through the use of Confidence Injection and Early Stopping.
-We fine-tune the model on this dataset using two standard learning paradigms: supervised fine-tuning (SFT), which directly imitates the revised reasoning chains, and Simple Preference Optimization (SimPO), which further amplifies the model’s preference for confident, efficient reasoning.
+To align LRM generation with the behavior encouraged by ConCISE, we first construct a fine-tuning dataset by applying the ConCISE pipeline to a set of questions.
+This yields concise reasoning chains in which redundant reflections are actively suppressed.
+We then fine-tune the LRM on this dataset using two standard learning paradigms: Supervised Fine-Tuning (SFT) and Simple Preference Optimization (SimPO).
 Both approaches optimize a shared objective that governs the desired generation behavior:
 $\begin{cases}\pi_{\theta}(S_{i})\rightarrow\pi_{\theta}(S_{i}+p_{i}),&\text{if }c_{i}<t_{i}\\ \pi_{\theta}(S_{i})\rightarrow\text{Terminate},&\text{if }i\geq i^{\prime}\text{ and }\hat{c}_{i}>t_{e}\end{cases}$
-This objective encourages the model to learn a generation policy guided by internal confidence, rather than merely imitating surface text patterns.
 Specifically, when the model’s confidence $c_{i}$ at a given step falls below the threshold $t_{i}$, it learns to approximate the distribution conditioned on an injected confidence phrase $p_{i}$, which helps improve internal certainty and suppress unnecessary reflections.
-After the first answer is generated (at step $i^{\prime}$), if the detected confidence score $\hat{c}_{i}$ exceeds the early stopping threshold $t_{e}$, the model is explicitly trained to terminate reasoning at that point.
-Through this process, the model learns to regulate its generation trajectory based on confidence, strengthening certainty during reasoning and halting once sufficient confidence is reached.
-As a result, the generation distribution favors precise and concise reasoning chains without compromising task performance.
+After the first answer is generated, if the detected confidence score $\hat{c}_{i}$ exceeds the early stopping threshold $t_{e}$, the model is explicitly trained to terminate reasoning at that point.
+Through fine-tuning, the model could learn to regulate its generation trajectory based on confidence, strengthening certainty during reasoning and halting once sufficient confidence is reached.
+(a) Average Acc, StepNum, and StepLen.
+(b) Analysis of Reflection steps.
+(c) Metrics of training datasets.
+Figure 5: Further analysis of reasoning chains and training datasets on DeepSeek-R1-Distill-Qwen-7B.
 4 Experiments
-Model Method Math-500 GSM8K AIME24 GPQA Average
-[…]
-DS-7B Origin 90.8 3854.2 100% 93.1 1441.9 100% 54.2 13573.5 100% 51.0 8141.8 100% 72.3 100%
-OverThinkSFT 92.2 2537.7 65.8% 93.0 1002.1 69.5% 52.5 11225.3 82.7% 51.1 7638.7 93.8% 72.2 78.0%
-SpiritSFT 91.0 2934.6 76.1% 91.4 1107.4 76.8% 51.7 11528.7 84.9% 50.3 7084.3 87.0% 71.1 81.2%
-ConCISESFT 92.0 2244.3 58.2% 92.9 831.9 57.7% 52.1 9750.8 71.8% 50.0 5891.5 72.4% 71.8 65.0%
-OverThinkSimPO 91.4 2405.1 62.4% 92.9 879.4 61.0% 50.0 9603.2 70.8% 49.9 6304.9 77.4% 71.0 67.9%
-SpiritSimPO 87.2 1765.1 45.8% 90.8 687.6 47.7% 38.3 6926.0 51.0% 50.2 5832.0 71.6% 66.6 54.0%
-ConCISESimPO 91.0 1945.7 50.5% 92.1 715.1 49.6% 48.3 7744.7 57.1% 48.0 4859.4 59.6% 70.0 54.2%
-DS-1.5B Origin 82.2 4783.7 100% 85.4 2219.0 100% 29.2 17464.9 100% 34.5 9492.3 100% 57.8 100%
-OverThinkSFT 84.0 3295.6 68.9% 85.2 1199.7 54.1% 30.0 12893.3 73.8% 34.8 8600.0 90.6% 58.5 71.8%
-SpiritSFT 83.8 3856.7 80.6% 84.8 1475.9 66.5% 28.8 13684.5 78.4% 32.3 8554.7 90.1% 57.4 78.9%
-ConCISESFT 83.6 2700.8 56.5% 84.9 923.3 41.6% 30.0 11359.0 65.0% 35.0 7252.6 76.4% 58.4 59.9%
-OverThinkSimPO 83.6 2737.8 57.2% 84.8 1002.9 45.2% 31.26 11465.1 65.7% 34.7 8249.8 86.9% 58.6 63.8%
-SpiritSimPO 82.6 2455.3 51.3% 82.6 804.2 36.2% 26.66 9946.3 57.0% 33.1 6910.3 72.8% 56.2 54.3%
-ConCISESimPO 83.6 2428.6 50.8% 84.3 803.0 36.2% 30.43 8810.1 50.4% 35.4 7056.1 74.3% 58.4 52.9%
-Table 1: Evaluation results of OverThink, SPIRIT, and ConCISE fine-tuned via SFT or SimPO on DeepSeek-R1-Distill 7B and 1.5B, showing accuracy(Acc.), token count(Tok.), and compression ratio(CR) across four benchmarks.
-Figure 4: Comparison of StepNum and StepLen across two models and four benchmarks.
-Stars indicate ConCISE-based methods, which achieve stronger stepwise compression than baselines while maintaining comparable accuracy.
 4.1 Settings
-Models.
-We evaluate ConCISE on two Large Reasoning Models (LRMs): DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-1.5B Guo et al.
-(2025).
-Although both models exhibit strong mathematical reasoning abilities, their substantial redundancy in reflection steps highlights the need for reasoning chain compression.
-Question Set.
-We find that effective reasoning-chain compression can be achieved with relatively limited data.
-Based on this observation, we select 2,000 questions from the MATH training setHendrycks et al.
-(2021) for fine-tuning, ensuring that each question yields a correct answer under greedy decoding and that the set maintains diversity in difficulty levels and response lengths.
-Detailed analysis is provided in the AppendixB.1.
-Evaluation Benchmarks.
+Models and Datasets.
+We evaluate ConCISE on four LRMs: DeepSeek-R1-Distill-Qwen-7B & 1.5BGuo et al.
+(2025), Skywork-OR1-7B-PreviewHe et al.
+(2025), and Qwen3-8BTeam (2025).
+As for the training dataset, we select 2,000 questions from the MATH training setHendrycks et al.
+(2021), ensuring that each question yields a correct answer under greedy decoding and maintaining diversity in difficulty and response lengths.
+Importantly, for each model, we generate its own fine-tuning data using the ConCISE pipeline, rather than relying on external models, which ensures consistency between the training data and the reasoning style of the target model.
+Evaluation.
 We evaluate model performance across four benchmarks: GSM8KCobbe et al.
 (2021), Math-500Hendrycks et al.
-(2021), AIME24Li et al.
-(2024), and GPQA_diamondRein et al.
+(2021), AIME24, and GPQA_diamondRein et al.
 (2024).
-All experiments use the same decoding configuration with temperature = 0.6 and top_p = 0.95Guo et al.
+All evaluations use the same decoding configuration with temperature = 0.6 and top_p = 0.95Guo et al.
 (2025).
 The maximum length is set to 16k for GSM8k, Math-500, GPQA_diamond, and 32k for AIME24.
-For AIME24 and GPQA_diamond, due to their higher difficulty and smaller sizes, we perform 8 sampling rounds per instance and report the mean values across all metrics to ensure stability and reliability.
-Evaluation Metrics.
-We adopt three primary metrics to comprehensively assess both reasoning accuracy and compression effectiveness.
+For AIME24 and GPQA_diamond, due to their higher difficulty and smaller sizes, we sample 8 times and report the mean values.
+As for metrics, we adopt three primary metrics to assess both reasoning accuracy and compression effectiveness comprehensively.
 Acc.
-denotes the final answer accuracy.
+denotes the accuracy of the final answer.
 Tok.
 refers to the average response length, measured in tokens.
-CR (Compression Rate) is defined as the ratio of the average response length to that of the original model, with lower values indicating higher compression.
-Baseline Methods.
-We compare our method against existing approaches that aim to remove redundant reasoning and construct efficient reasoning datasets.
-OverThink.Chen et al.
+CR (Compression Rate) is defined as the ratio of the average response length to that of the original model, with lower values indicating better compression.
+Baselines.
+We compare our method against existing approaches that aim to remove redundant reasoning and construct efficient reasoning datasets, with further comparisons after fine-tuning under SFT and SimPO scenarios.
+OverThink Chen et al.
 (2024) samples eight responses from the LRM with a relatively high temperature and selects the shortest correct one.
 Only the first reflection step after the answer is retained.
-Spirit.Cui et al.
+Spirit Cui et al.
 (2025) iteratively removes steps with minimal impact on perplexity (PPL), deleting 30% of the least critical steps.
-Adjacent contexts are merged when needed to maintain coherence.
-Training Details.
-We construct training datasets based on the question set using ConCISE, OverThink, and Spirit.
-For Spirit, compression is applied to reasoning chains generated via greedy search.
-During training, we adopt two strategies: SFT and SimPO.
-SFT directly uses the question and its corresponding compressed reasoning trace for supervised fine-tuning.
-For SimPO, reject samples are also required Rafailov et al.
-(2023).
-Following the original OverThink setup, we use the longest correct reasoning chain among the eight samples as the reject samples.
-To ensure a fair comparison, ConCISE adopts the same strategy.
-For Spirit, the full reasoning chain before compression serves as reject samples.
-All models are trained with $lr=1e-6$; SFT runs for 2 epochs and SimPO for 1 epoch.
+Adjacent contexts are merged to maintain coherence.
+Appendix B.2 shows more implementation and training details.
 4.2 Reults and Analysis
-This section evaluates multiple compression methods across various reasoning benchmarks.
+This section evaluates different compression methods across various reasoning benchmarks.
 Our results show that ConCISE strikes a superior balance between compression and task performance compared to baseline methods, enabling models to effectively eliminate redundant reasoning steps.
 Additionally, models fine-tuned with ConCISE on mathematical datasets generalize well to GPQA_diamond, showcasing its robustness.
-4.2.1 ConCISE achieves superior compression-performance balance
-As shown in Table 1, ConCISE demonstrates strong reasoning chain compression under both SFT and SimPO settings while maintaining comparable accuracy.
-Under SFT, it achieves a compression rate of ~60%, and under SimPO, it further improves compression to ~50%.
-In contrast, OverThink achieves competitive task performance but exhibits weaker compression under SFT.
-Although it achieves further compression under SimPO, its final compression rate and task performance are comparable to, or even slightly worse than, ConCISE under SFT.
-As for Spirit, it underperforms in both settings, achieving the weakest compression under SFT and suffering noticeable task performance degradation compared to the other two methods.
-While Spirit matches ConCISE’s compression rate under SimPO, its accuracy drops more severely, especially on challenging datasets like AIME24.
-Overall, ConCISE offers a better balance between compression and performance compared to baseline methods.
-(a) Step-level analysis of reasoning chains on Math-500
-(b) Comparison of training datasets of different methods
-Figure 5: Extended analysis of reasoning chain characteristics. More details can be found in the Appendix B.3.
-4.2.2 How does ConCISE achieve better compression-performance balance?
-To deeply understand how ConCISE achieves a better balance between compression and performance, we perform a stepwise analysis of generated reasoning chains on benchmarks.
-As shown in Figure 4, we evaluate the average number of reasoning steps in the thinking stage(StepNum) and the average token count per step in the thinking stage(StepLen) across different benchmarks.
-ConCISE consistently achieves the fewest reasoning steps, while StepLen remains similar to that of the original model.
-Further details can be found in the AppendixA.3.
-Further analysis, focusing on reflection steps in the generated reasoning chains on MATH-500, reveals how each method handles reflection behavior.
-We use several metrics, including StepNum, Non-RefNum, RefNum, Pre-FAS RefNum, and Post-FAS RefNum.
-Non-RefNum represents the number of non-reflection steps, and the last three metrics measure the number of reflection steps across the entire reasoning chain, before FAS(First Answer Step), and after FAS, respectively.
-The results are shown in Figure 5(a).
-We find that, in terms of RefNum, Spirit-SFT retains the most reflection steps, followed by OverThink, while ConCISE and Spirit-SimPO retain fewer reflection steps.
-For Non-RefNum, OverThink and ConCISE perform similarly, while Spirit-SimPO slightly underperforms and also causes a performance drop.
-This suggests that none of the compression methods can effectively reduce non-reflection steps without compromising model performance.
-Consequently, it indicates that existing compression methods mainly achieve compression by pruning reflection steps.
-For Pre-FAS RefNum, OverThink scores the highest, followed by Spirit, as neither method effectively controls early-stage reasoning.
-Spirit-SFT, on the other hand, has the highest Post-FAS RefNum, indicating its limited ability to identify and prune redundant steps.
-In contrast, ConCISE maintains consistent control over both pre- and post-answer reflections.
-These differences help explain the performance gaps: OverThink’s inability to prune pre-answer reflections limits its compression potential, while Spirit’s failure to identify redundant steps leads to suboptimal pruning, removing essential content and degrading performance.
-ConCISE, through confidence injection and early stopping, precisely removes redundant reflection steps without harming essential reasoning process.
-4.2.3 How does the training dataset affect compression?
-Since all methods share identical training configurations and hyperparameters, the quality of the training dataset directly determines the final compression performance.
-To further investigate which features of the training dataset are highly correlated with the final compression performance, we analyzed the reasoning datasets constructed by ConCISE, OverThink, Spirit, Greedy-Search (reject samples of Spirit), and Longest-of-N (reject samples of ConCISE and OverThink), with the evaluation metrics including Tok., StepNum, StepLen, RefNum, Pre-FAS RefNum, and Post-FAS RefNum.
-Figure 5(b) shows the distribution of these metrics across the training datasets.
-Our results indicate that surface-level metrics such as Tok., StepNum, and StepLen do not directly correlate with the compression rate.
-For instance, ConCISE and OverThink have nearly identical values on these metrics, but ConCISE achieves much better compression after fine-tuning.
-In contrast, other metrics reveal clear distinctions: OverThink records the highest Pre-FAS RefNum, while Spirit leads in Post-FAS RefNum.
-ConCISE, by comparison, consistently maintains lower values on both.
-This aligns with the reasoning chain analysis in the previous subsection, indicating that key features from the training data are effectively transferred to the fine-tuned model.
-These features enable the model to compress reasoning, particularly reflection steps, in line with the dataset design principles.
-4.2.4 Can reasoning compression methods generalize across tasks?
-Another important question is whether compression methods can generalize beyond the training data distribution.
-This property is crucial for the practical use of reasoning chain compression.Qu et al.
-(2025) In our setup, the training dataset is constructed from the question set of the MATH training set.
-Accordingly, MATH-500 serves as the in-distribution evaluation, GSM8K and AIME24 represent in-domain but out-of-distribution tasks, and GPQA_diamond serves as an out-of-domain task, differing substantially in content, format, and prompt structure.
-As shown in Table 1, ConCISE consistently achieves a better balance between compression and performance than baseline methods across both in-domain and out-of-domain tasks.
-These results indicate that fine-tuning on efficient reasoning chains constructed by ConCISE enables models to structurally suppress redundant reflection steps, a property that holds regardless of task type or prompt format, without compromising overall performance.
+ConCISE achieves superior balance between compression and task performance
+As shown in Table 1, ConCISE consistently demonstrates strong compression ability across all four LRMs and under both SFT and SimPO settings, while maintaining model performance.
+Particularly under SimPO settings, ConCISE achieves a compression rate of ~50% with minimal impact on the original model’s performance, achieving excellent compression results on the challenging task AIME24 as well as the out-of-domain task GPQA.
+In contrast, OverThink achieves competitive task performance but exhibits weaker compression under both SFT and SimPO settings.
+Regarding Spirit, despite attaining a compression rate comparable to that of ConCISE under SimPO settings, its task performance suffered a notable degradation, especially on challenging tasks like AIME24.
+Overall, compared to baseline methods, ConCISE offers a better balance between compression and performance, demonstrating excellent in-domain and out-of-domain generalization and robustness.
+How does ConCISE achieve a better compression & performance balance?
+To better understand how ConCISE achieves an improved balance between compression and performance, we particularly analyze the structure of reasoning chains generated on various benchmarks of DeepSeek-R1-Distill-Qwen-7B.
+As shown in Figure 5(a), we evaluate the average number of steps (StepNum) and the average token count per step (StepLen).
+ConCISE generates the fewest reasoning steps, while its StepLen remains comparable to that of the original model.
+More detailed analysis focuses on reflection steps within reasoning chains on MATH-500, revealing distinct compression strategies among the methods.
+As shown in Figure 5(b), we find that pruning reflection steps is key to compression, and preserving non-reflection steps is paramount for model performance.
+ConCISE uniquely achieves this critical balance.
+Specifically, Spirit-SFT retains the most reflection steps (RefNum), followed by OverThink, while ConCISE retains the fewest.
+Regarding non-reflection steps (Non-RefNum), OverThink and ConCISE exhibit similar counts.
+Spirit shows slightly fewer non-reflection steps under SimPO, albeit at the cost of a performance drop.
+This suggests that current compression techniques primarily achieve compression by pruning reflection steps, as none of the evaluated methods can effectively reduce non-reflection steps without impairing model performance.
+Further dissecting the reflection steps, OverThink and Spirit have the most reflection steps before FAS(Pre-FAS RefNum) due to ineffective early-stage reasoning control.
+In contrast, Spirit-SFT has most reflection steps after FAS(Post-FAS RefNum), indicating its inability to prune redundant post-answer reflections.
+These distinct strategies in reflection management contribute to performance disparities: OverThink’s limited pre-answer pruning curtails its compression potential, while Spirit’s difficulty in accurately identifying redundant steps often degrades performance by removing essential content.
+ConCISE, however, achieves a better compression-performance balance by precisely eliminating redundant reflection steps throughout the entire reasoning chain without compromising the essential reasoning process.
+We present more details about the above analysis in Appendix B
+How does the training dataset affect compression?
+Since all methods operate under identical training configurations, variations in final compression performance are fundamentally rooted in the characteristics of their respective training datasets.
+To this end, we analyze training datasets constructed by ConCISE, OverThink, Spirit, Greedy-Search (derived from Spirit’s rejected samples), and Longest-of-N (derived from rejected samples of ConCISE and OverThink).
+As shown in Figure 5(c), surface-level metrics such as response tokens (Tok.), step count (StepNum), and average step length (StepLen) show no strong direct correlation with the final compression rate.
+For instance, while the training datasets of ConCISE and OverThink yield nearly identical values for these metrics, ConCISE demonstrates significantly better compression after fine-tuning under both SFT and SimPO.
+Conversely, Metrics related to reflection offer more insightful distinctions among the methods.
+For instance, OverThink has the highest values for Pre-FAS RefNum, while Spirit shows the highest Post-FAS RefNum.
+ConCISE, in contrast, consistently maintains lower figures for both these metrics, underscoring its more effective control over the model’s reflective tendencies.
+Echoing the analysis from the previous subsection, these observed characteristics within the training datasets directly mirror the fine-tuned model’s capacity for reflection management.
+This provides strong evidence that the model effectively learns features about reflection from the training data during the fine-tuning process.
+Thus, it becomes clear how ConCISE uses its unique training data design to achieve significant compression by identifying and cutting down on unnecessary reflection steps.
+4.3 Ablation Study
 Benchmark Method Acc. Tok. StepNum StepLen CR
-Math-500 Origin 90.8 3854.2 113.7 33.9 100%
-ConCISE 92.0 2244.3 63.0 35.6 58.2%
-Conf. Inj. 91.8 2619.6 79.5 33.0 68.0%
-Early Stop 92.4 2653.8 68.4 38.8 68.9%
-GSM8K Origin 93.1 1441.9 41.4 34.8 100%
-ConCISE 92.9 831.9 22.5 37.1 57.7%
-Conf. Inj. 92.6 933.8 23.2 40.2 64.8%
-Early Stop 92.7 1003.0 22.3 45.0 69.6%
-AIME24 Origin 54.2 13573.5 463.0 29.3 100%
-ConCISE 52.1 9750.8 333.0 29.3 71.8%
-Conf. Inj. 51.3 10165.8 372.2 27.3 74.5%
-Early Stop 54.2 12204.6 389.7 31.3 89.9%
-Table 2: Ablation study of DeepSeek-R1-Distill-Qwen-7B across three mathematical benchmarks.
-4.3 Ablation Study
+Math-500 Origin 90.8 3854 113.7 33.9 100%
+ConCISE 92.0 2244.3 63.0 35.6 58%
+Conf. Inj. 91.8 2620 79.5 33.0 68%
+Early Stop 92.4 2654 68.4 38.8 69%
+GSM8K Origin 93.1 1442 41.4 34.8 100%
+ConCISE 92.9 832 22.5 37.1 58%
+Conf. Inj. 92.6 934 23.2 40.2 65%
+Early Stop 92.7 1003 22.3 45.0 70%
+AIME24 Origin 54.2 13574 463.0 29.3 100%
+ConCISE 52.1 9751 333.0 29.3 72%
+Conf. Inj. 51.3 10166 372.2 27.3 75%
+Early Stop 54.2 12205 389.7 31.3 90%
+Table 2: Ablation study of Confidence Injection and Early Stopping on DeepSeek-R1-Distill-Qwen-7B.
 We conduct an ablation study to evaluate the two core components of ConCISE: Confidence Injection and Early Stopping.
-For each mechanism, we construct a training dataset and fine-tune the base model using the same configurations as in the main experiments.
-All experiments are performed in the supervised fine-tuning (SFT) setting and evaluated across three mathematical reasoning tasks: Math-500, GSM8K, and AIME24.
+For each, we construct a training dataset and fine-tune DeepSeek-7B using the same configurations as in the main experiments.
+All experiments are performed under SFT settings and evaluated across three mathematical reasoning tasks: Math-500, GSM8K, and AIME24.
 The results, summarized in Table 2, show that both mechanisms achieve similar reasoning accuracy to full ConCISE but exhibit weaker compression performance.
-Both methods focus on reducing redundant reflection steps while maintaining the integrity of the reasoning process.
-Confidence Injection suppresses unnecessary iterations by reinforcing the model’s confidence, but it has a limited impact on terminating reasoning after the final answer.
+Specifically, Confidence Injection suppresses unnecessary reflections by boosting the model’s confidence, but has a limited impact on terminating reasoning after the final answer.
 Early Stopping, on the other hand, halts excessive post-answer reasoning but does not address earlier stages.
-Interestingly, Early Stopping nearly preserves the original model’s accuracy on benchmarks like AIME24, suggesting that post-FAS reflection steps often contain substantial redundancy and can be safely stopped without harming task performance.
-In summary, Confidence Injection reduces redundancy across the entire reasoning sequence, while Early Stopping targets post-answer overthinking.
-Only by combining both can ConCISE achieve comprehensive redundancy elimination, resulting in superior compression without compromising reasoning quality.
+Only by combining both can we achieve comprehensive redundancy elimination without compromising reasoning quality.
+Figure 6: Average performance of training-free methods and ConCISE variants on DeepSeek-R1-Distill-Qwen-7B and 1.5B across GSM8K, Math-500, and AIME24.
+4.4 Analysis of Training-free Methods
+To further demonstrate the advantages of the ConCISE framework, we compare it against representative training-free methods.
+Specifically, we include TALE Han et al.
+(2024), which constrains reasoning length by injecting explicit token budgets into prompts, and DEER Yang et al.
+(2025), which dynamically decides when to terminate reasoning by monitoring the generation probability of the answer.
+For DEER, we reproduce two variants with $\lambda=0.95$ and $\lambda=0.98$.
+We also implement a training-free version of our method, denoted ConCISE-Decoding, which directly applies Algorithm 1during inference to decide when to terminate reasoning, without relying on ground-truth verification.
+We evaluate these methods on GSM8K, Math-500, and AIME24, and report the averaged results in Figure 6, with detailed statistics provided in Table 6.
+We find that training-free methods can achieve partial compression, but their compression levels are unstable and often come with substantial accuracy degradation.
+For instance, DEER achieves relatively stable compression but introduces a 5–8% accuracy drop.
+Meanwhile, TALE attains a high compression ratio on GSM8K with DeepSeek-7B, but performs poorly on other datasets and models.
+In contrast, ConCISE-SimPO consistently achieves a superior overall trade-off, and even the training-free variant ConCISE-Decoding outperforms other training-free approaches.
+These results highlight that ConCISE enables more precise elimination of redundant reflection steps, thereby yielding higher-quality reasoning chains.
 5 Conclusion
-In this work, we first modeled the generation of reflection steps in Large Reasoning Models (LRMs) from a confidence perspective.
-This modeling approach enabled us to systematically identify two primary patterns of redundancy: Confidence Deficit, where models unnecessarily undertrust intermediate steps, and Termination Delay, where reflection persists despite reaching a confident answer.
-To address these issues, we introduced ConCISE, a novel confidence-guided framework designed to enhance reasoning efficiency.
-ConCISE employs two targeted mechanisms: Confidence Injection to suppress redundant reflections by boosting internal confidence, and Early Stopping to halt generation upon detecting sufficient answer confidence.
-Our experiments empirically validate that ConCISE significantly compresses response length (~50% under SimPO) while preserving strong accuracy across diverse reasoning benchmarks.
-Notably, ConCISE achieves superior compression-accuracy trade-offs over existing methods and exhibits robust generalization to out-of-domain tasks.
-This study offers a principled approach to modeling and mitigating redundancy in LRM reasoning, enabling more resource-efficient and user-friendly AI systems.
+In this work, we first modeled the generation of reflection steps in LRMs from a confidence-guided perspective, helping to understand two key patterns responsible for redundant reflection within the reasoning process of LRMs: Confidence Deficit and Termination Delay.
+Based on this, we propose ConCISE, which employs Confidence Injection and Early Stopping mechanisms to respectively address the above two patterns to suppress redundant reflection steps and produce efficient, concise reasoning chains.
+Extensive experiments demonstrate that LRMs fine-tuned on ConCISE-generated data significantly compress their responses while maintaining strong task performance.
 Limitations
 While ConCISE demonstrates strong performance in compressing reasoning chains, it is subject to certain limitations that merit further exploration:
-Further Compression of Step Length.
-While ConCISE effectively targets redundant reflection steps, analysis indicates substantial potential remains for intra-step compression based on average step length.
-Future work could explore integrating techniques like TokenSkip with ConCISE for greater compression.
+Room for Further Compression.
+While ConCISE effectively addresses redundant reflection steps, analysis indicates that significant potential remains for further compression of both non-reflection steps and the average length of each step.
+Future work could explore integrating techniques with ConCISE to achieve a higher degree of compression while maintaining model performance.
 Confidence Estimation During Reasoning.
-Although ConCISE introduces a lightweight confidence detector for the post-answer phase, it currently lacks a direct mechanism to model confidence prior to the first answer, relying instead on reflection occurrences as an indirect proxy.
-Model Dependency and Sensitivity.
-Validated on DeepSeek models, ConCISE’s effectiveness and the optimal tuning of its components (e.g., confidence phrases, detector, thresholds) may be sensitive to model architecture or scale.
-Its performance across diverse LRMs requires further study, potentially necessitating model-specific adaptations.
+Although ConCISE introduces a lightweight confidence detector for the post-answer phase, it still lacks a direct mechanism to model confidence before the first answer.
+Instead, it relies on the occurrence of reflection steps as an indirect proxy.
+Future work could explore training a lightweight model to detect the model’s internal confidence, thereby enabling more fine-grained control.
+Integration with RLVR methods.
+This work focuses on constructing concise reasoning traces and validating them under SFT and SimPO, which we consider sufficient to support our core contribution.
+At the same time, we recognize that RLVR methods such as GRPOShao et al.
+(2024) are powerful and widely used in post-training.
+Exploring how confidence-guided signals can be incorporated into RLVR setups is a valuable future direction.
+Acknowledgement
+This research was supported in part by the National Natural Science Foundation of China under Grant No.
+62402267, 62432004, the China Postdoctoral Science Foundation under Grant No.
+BX20240177, 2025M771499, and a grant from the Guoqiang Institute, Tsinghua University.
 Appendix A Details about ConCISE
@@
 To mitigate potential overfitting to any specific phrase, one phrase is randomly selected from the pool for injection at runtime.
+Additionally, another crucial mechanism within the Confidence Injection process is identifying whether the current step constitutes a reflection step.
+The accurate and swift identification of such steps is paramount for the success of Confidence Injection.
+Consequently, informed by our observations of reflection steps during LRM reasoning processes, we employed a rule-based detection method.
+This method, which is detailed in Appendix A.3.1, achieves reasonably accurate detection while having a negligible impact on the model’s inference speed.
 Table 3: The Pool of 20 Confidence Phrases used for Confidence Injection.
@@
 The effectiveness of the Early Stopping mechanism relies heavily on the accuracy of its confidence detector.
-To identify an appropriate probing prompt, we analyzed the reasoning chains generated by the DeepSeek-R1-Distill-Qwen-7B model using Greedy decoding on the training question set.
-This analysis revealed that the model frequently expresses confidence during the later part of the reasoning chain.
-Table 4 summarizes the frequencies of the most common self-reported confidence expressions observed within chains.
-Table 4: Frequency of Common Confidence Expressions observed in 2000 Reasoning Chains.
-Confidence Expression Frequency Count
-I’m confident 423
-I feel confident 170
-I can be confident that 157
-I can confidently say 70
-This gives me confidence that 68
-I’m pretty confident 43
-I’m pretty sure 37
-Based on the prevalence of phrases starting with "I’m" leading into confidence statements, we selected "So, I’m" as the fixed probing prompt.
-Subsequently, we identified "confident", "pretty confident", "sure", and "pretty sure" as the primary confidence-indicative tokens.
+To identify an appropriate probing prompt, we first analyzed the reasoning chains generated by four Large Reasoning Models (LRMs) using Greedy decoding on the training question set.
+This analysis revealed that the model frequently expresses confidence during the later part of the reasoning chain, and as shown in Table 4, self-reported confidence expressions starting with “I’m” (such as “I’m confident”) are very common.
+Based on this observation, particularly considering the prevalence of confidence statements led by “I’m”, we considered "So, I’m" as a potential probing prompt.
+Concurrently, we also experimented with alternative probing prompts, such as "So, I feel" and "I can be".
+However, we observed that these alternative prompts often resulted in high confidence scores before the model had reached the final answer (i.e., prior to Final Answer Serialization, FAS), leading to unnecessary computational overhead.
+Therefore, considering these factors—namely, the prevalence of “I’m”-led confidence statements in the model’s natural expressions and the shortcomings of the alternative prompts—we ultimately selected "So, I’m" as the fixed probing prompt.
+Subsequently, we identified the primary confidence-indicative tokens that follow the "So, I’m" prompt, which are: "confident", "pretty confident", "sure", and "pretty sure".
+Table 4: Frequency of Common Confidence Expressions Observed in 2000 training Reasoning Chains Across Different Large Reasoning Model Series(DeepSeek, Qwen and Skywork)
+Expression DeepSeek-7B DeepSeek-1.5B Qwen3-8B Skywork-OR17B
+I’m confident 423 335 127 560
+I feel confident 170 172 58 144
+I can be confident that 157 156 475 130
+I can confidently say 70 176 91 98
+This gives me (more)confidence that 68 81 23 86
+I’m pretty confident 43 69 2 60
+I’m pretty sure 37 28 2 17
+Therefore, confident that 1 0 26 5
+Total (sum of listed expressions) 969 1017 804 1100
 The confidence score, $\hat{c}_{i}$, for a given state $S_{i}$ is then calculated by summing the probabilities of generating these tokens or sequences immediately following the probing prompt "So, I’m":
-[…]
-[…]
-[…]
-[…]
-[…]
+$\hat{c}_{i}=$ $P(\text{"confident"}\mid S_{i},\text{"So, I'm"})$
+$+P(\text{"sure"}\mid S_{i},\text{"So, I'm"})$
+$+P(\text{"pretty"}\mid S_{i},\text{"So, I'm"})\times$
+$\big(P(\text{"confident"}\mid S_{i},\text{"So, I'm pretty"})$
+$\quad+P(\text{"sure"}\mid S_{i},\text{"So, I'm pretty"})\big)$
 Here, notation like $P(\text{token}\mid S_{i},\text{sequence})$ represents the conditional probability of generating the next token given the preceding state $S_{i}$ and the specific text sequence provided as context.
@@
   • Difficult problems or those with less rigorous reasoning undergo two or more rounds of reflection.
-To find a threshold value that aligns with this principle, we empirically evaluated $t_{e}\in\{0.4,0.5,0.6,0.7\}$.
-For each value, we measured the probability distribution of the stopping points across different stages: stopping after the initial answer generation (0 reflections), after the first reflection, after the second reflection, etc.
+To find a threshold value that aligns with this principle, we empirically evaluated $t_{e}\in\{0.4,0.5,0.6,0.7\}$ on the DeepSeek-R1-Distill-Qwen-7B model.
+For each value, we measured the probability distribution of the exceeding points(when $\hat{c}_{i}>t_{e}$) across different stages: exceeding after the initial answer generation (0 reflections), after the first reflection, after the second reflection, etc.
 The results of this analysis are illustrated in Figure 7.
-Figure 6: Distribution of early stopping points (probability of stopping at Pre-FAS, FAS, Ref1, 2, …, 5) for different threshold values ($t_{e}$).
-Observing the distributions in Figure 6, we found that setting the threshold $t_{e}=0.5$ yielded the stopping behavior that best matched our desired criteria across varying problem difficulties.
-Therefore, we selected $t_{e}=0.5$ as the definitive early stopping threshold for all subsequent experiments conducted in this work.
+Figure 7: Distribution of exceeding points(when $\hat{c}_{i}>t_{e}$) probability of Pre-FAS, FAS, Ref1, 2, …, 5 for different threshold values ($t_{e}$).
+Benchmark $T_{e}$ Acc. Tok. StepNum StepLen CR
+Math-500 0.4 91.4 2390.6 68.4 34.9 62.0%
+0.5 92.0 2244.3 63.0 35.6 58.2%
+0.6 91.8 2324.3 66.1 35.2 60.3%
+0.7 92.0 2345.1 67.5 34.8 60.8%
+GSM8K 0.4 92.4 828.9 19.9 41.6 57.5%
+0.5 92.9 831.9 22.5 37.1 57.7%
+0.6 93.0 831.7 20.4 40.7 57.7%
+0.7 92.7 849.3 21.7 39.2 58.9%
+AIME24 0.4 45.8 11548.7 402.1 28.7 85.1%
+0.5 52.1 9750.8 333.0 29.3 71.8%
+0.6 52.1 9802.5 340.5 28.8 72.2%
+0.7 51.7 9719.9 339.8 28.6 71.6%
+Table 5: Effect of different early stopping thresholds (0.4, 0.5, 0.6, 0.7) on the SFT performance of DeepSeek-R1-Distill-Qwen-7B across mathematical benchmarks.
+Observing the distributions in Figure 7, we found that these $t_{e}$ settings generally cause the model to achieve a confidence score $\hat{c}_{i}>t_{e}$ at the position of the first reflection, which aligns with our three guiding principles.
+Furthermore, the confidence monitored by our designed lightweight detector rarely indicates high confidence before the FAS (Final Answer Serialization) step.
+This is advantageous as it effectively reduces the computational overhead associated with prematurely judging the correctness of an answer, making it highly compatible with the ConCISE methodology.
+Therefore, we further conducted experiments for $t_{e}\in\{0.4,0.5,0.6,0.7\}$ under the SFT setting of DeepSeek-R1-Distill-Qwen-7B, and Table 5 presents these experimental results.
+We found that under the settings of $t_{e}=0.5,0.6,$ and $0.7$, the model ultimately exhibited similar performance in both compression rate (CR) and accuracy (Acc), demonstrating robustness to the choice of $t_{e}$.
+Consequently, we selected $t_{e}=0.5$ for subsequent extensive experiments.
+However, when $t_{e}=0.4$, there was a significant decline in performance.
+This is because when $t_{e}$ is set to $0.4$, more data exists at the FAS step, preventing the model from adequately checking its reasoning process; this excessive confidence leads to a performance drop.
 A.3 Reflection step detection
@@
 This method checks for the presence of predefined "reflection keywords" within the generated text of a given step, $S_{i}$.
-[…]
-[…]
+Formally, let $K$ be the predefined set of reflection keywords.
+A step $s_{i}$ is identified as the start of a reflection, denoted as $\text{IsReflectionStart}(s_{i})$, if its text contains any keyword $k$ from the set $K$:
+$\text{IsReflectionStart}(s_{i})\equiv\bigvee_{k\in K}(k\subseteq s_{i})$
 where $k\subseteq s_{i}$ indicates that the keyword $k$ appears as a substring of step $s_{i}$.The set $K$ of reflection keywords used for rule-based detection consists of the following:
-  • "wait", "alternatively", "check", "reconsider", "reflect", "rethink", "reconsidering", "reviewing", "reassess", "pause", "second thought", "reevaluate", "verify".
+  • "wait", "alternatively", "check", "reconsider", "reflect", "rethink", "reconsidering", "reviewing", "reassess", "pause", "second thought", "reevaluate", "verify", "think again".
 While effective for identifying the onset of a reflection sequence, this method typically cannot recognize subsequent steps within the same reflection process, as these later reflection steps often lack the initial triggering keywords.
@@
 Similarly, this rule-based detection is used for the Figure 4(a) analysis to determine if an intervention (like inserting a confidence phrase) triggers the start of a new reflection, as only detecting the initial step with this lightweight method is necessary.
+(a) Difficulty diversity of question set.
+(b) Length diversity of question set.
+Figure 8: Diversity of the question set, showing distributions for difficulty and length.
 A.3.2 LLM-as-a-Judge
@@
 Consequently, for the implementation of the OverThink baseline comparison, reasoning chains longer than 3000 tokens were manually annotated to ensure accuracy.
-Furthermore, the LLM-as-a-Judge approach is the basis for calculating several reflection-based metrics presented in our results, such as RefNum, Non-RefNum, pre-FAS RefNum, and post-FAS RefNum (shown in Figure 5(a) and Figure 5(b)).
+Furthermore, the LLM-as-a-Judge approach is the basis for calculating several reflection-based metrics presented in our results, such as RefNum, Non-RefNum, pre-FAS RefNum, and post-FAS RefNum (shown in Figure 5(b) and Figure 5(c)).
+Model Method GSM8K Math-500 AIME24 Average
+Acc. CR Acc. CR Acc. CR Acc. CR
+DeepSeek-7B Origin 93.1 100% 90.8 100% 54.2 100% 79.4 100%
+DEER(0.98) 90.3 59% 87.0 51% 47.1 69% 74.8 60%
+DEER(0.95) 87.9 56% 85.2 49% 43.4 64% 72.2 56%
+TALE 89.0 47% 90.0 95% 52.9 98% 77.3 80%
+ConCISE-Decoding 92.7 62% 89.2 55% 49.6 67% 77.2 61%
+ConCISE-SimPO 92.1 50% 91.0 51% 48.3 57% 77.1 53%
+DeepSeek-1.5B Origin 85.4 100% 82.2 100% 29.2 100% 65.6 100%
+DEER(0.98) 79.7 41% 77.6 54% 24.6 70% 60.6 55%
+DEER(0.95) 74.7 43% 76.4 49% 21.3 64% 57.5 52%
+TALE 84.9 90% 81.6 98% 29.6 98% 65.4 95%
+ConCISE-Decoding 83.9 53% 79.0 52% 26.3 52% 63.1 52%
+ConCISE-SimPO 84.3 36% 83.6 51% 30.4 50% 66.1 46%
+Table 6: Detailed results of training-free baselines and ConCISE variants on DeepSeek-R1-Distill-Qwen-7B and 1.5B across GSM8K, Math-500, and AIME24.
+Acc: accuracy (%), CR: compression ratio relative to original (lower CR indicates shorter outputs).
+(a) Average Acc, StepNum, and StepLen.
+(b) Analysis of Reflection steps.
+(c) Metrics of training datasets.
+Figure 9: Further analysis of reasoning chain and training dataset characteristics on DeepSeek-R1-Distill-Qwen-1.5B.
 Appendix B Details about Experiments
-B.1 Question Set
-Figure 7: Difficulty diversity of question set.
-Figure 8: Length diversity of question set.
-While the construction methodology for the question set is detailed in Section 4, here we focus on illustrating its diversity in terms of difficulty and length.Figures 7 and 8 demonstrate the diversity of the Question set.
+B.1 Datasets
+Question Set
+While the construction methodology for the question set is detailed in Section 4, here we focus on illustrating its diversity in terms of difficulty and length.Figures 8(a) and 8(b) demonstrate the diversity of the Question set.
 Difficulty distribution is derived from metric Level in the MATH dataset, while length distribution reflects reasoning chain lengths from DeepSeek-R1-Distill-Qwen-7B under greedy decoding.
-B.2 Validation Dataset
+Furthermore, it is important to note a subsequent refinement process for these selected questions when constructing the final training set.
+Although the initial 2000 data entries were chosen based on the premise that the model could correctly answer them using greedy search, the generation processes for our ConCISE and OverThink methods do not strictly adhere to greedy decoding, and ConCISE will discard wrong reasoning chains.
+Therefore, for training purposes, we ultimately utilize the common subset of correctly reasoned chains derived from the outputs of three methods: ConCISE, OverThink, and Spirit.
+This intersection results in a refined training dataset of approximately 1900 samples.
+Validation Dataset
 For the experiments presented in Figure 4, we curated the Verification Dataset.
@@
 This curated set facilitates reliable LLM-based annotation for the validation experiments.
-B.3 Evaluation Dataset
-The analysis presented in Figure 5 required LLM-as-a-Judge annotations.
+Evaluation Dataset
+The analysis presented in Figure 5(b) and Figure 9(b) required LLM-as-a-Judge annotations.
 To ensure reliable results despite this method’s limitations with long contexts, we filtered the MATH500 dataset.
@@
 This yielded a common subset of 340 problems, on which our subsequent comparative analysis of the different methods was exclusively performed.
+B.2 Training Details
+Models.
+DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-1.5B are reasoning models derived by DeepSeek from the Qwen2.5 series models Yang et al.
+(2024)through distillation training, using training and generation data from the DeepSeek-R1 model.
+Skywork-OR1-7B-Preview was further developed via reinforcement learning, exhibiting strong capabilities on test sets such as mathematics.
+Meanwhile, Qwen3-8B is a recently released hybrid reasoning model noted for its powerful reasoning abilities.
+All four of these are open-source models.
+While they possess strong Chain-of-Thought (CoT) reasoning capabilities, they also present clear issues with reasoning redundancy.
+Consequently, we employ these four models for subsequent training and evaluation.
+Furthermore, we ensure that these models, along with all datasets involved in the training and evaluation processes, are utilized following their original licenses and intended purposes.
+Training.
+We construct training datasets based on the question set using ConCISE, OverThink, and Spirit.
+For the generation process of ConCISE, we set Temperature=0.6, Top_p=0.95.
+For OverThink, we sample 8 reasoning chains under Tempature=1.0(which is relatively high due to the recommended temperature for solving math problems of the LRMs is usually 0.6-0.8), Top_p=0.95.
+For Spirit, compression is applied to reasoning chains generated via greedy search.
+During training, we adopt two strategies: SFT and SimPO.
+SFT directly uses the question and its corresponding compressed reasoning trace for supervised fine-tuning.
+For SimPO, reject samples are also required Rafailov et al.
+(2023).
+Following the original OverThink setup, we use the longest correct reasoning chain among the eight samples as the reject samples.
+To ensure a fair comparison, ConCISE adopts the same strategy.
+For Spirit, the full reasoning chain before compression serves as reject samples, because this is precisely the comparative effect Spirit aims to demonstrate: important steps are retained while preference relationships are learned only for non-important steps.
+We use LlamaFactory for both SFT and SimPO fine-tuning Zheng et al.
+(2024).
+All models are trained with lr=1e6lr=1e-6; SFT runs for 2 epochs and SimPO for 1 epoch.
+The effective batch size for all training is 32 (using per_device_train_batch_size=1 and gradient_accumulation_steps=8 on 4 A800 GPUs), and max_length is set to 10240.
+B.3 Analysis on DeepSeek-1.5B
+Regarding the analysis conducted on DeepSeek-R1-Distill-Qwen-7B in the Experimental Section 4.1, we also performed a similar analysis for DeepSeek-R1-Distill-Qwen-1.5B.
+Figure 9 shows the specific results.
+The trends exhibited by the data, as well as the relationships among the three methods, are largely consistent with those presented for DeepSeek-R1-Distill-Qwen-7B.
+Consequently, the final conclusions are also consistent.
 Prompt You are an AI assistant trained to analyze reasoning steps in a response.
@@
 Step6: Let me count: Tuesday plus one day is Wednesday, plus two days is Thursday, plus three days is Friday, plus four days is Saturday, and plus five days is Sunday.
-So, 5 days after Tuesday is $𝑆𝑢𝑛𝑑𝑎𝑦$.
+So, 5 days after Tuesday is Sunday.
 Step7: Therefore, 40 days from Tuesday is Sunday.
@@
 Yeah, that seems right.
-Step8: Final Answer: The day of the week 40 days from Tuesday is $𝑆𝑢𝑛𝑑𝑎𝑦$.
+Step8: Final Answer: The day of the week 40 days from Tuesday is Sunday.
 Figure 11: An example input of using the LLM-as-a-Judge method for reflection step detection.